### v1

# MaskMoE: Boosting Token-Level Learning via Routing Mask in Mixture-of-Experts

Abstract
Scaling model capacity enhances its capabilities but significantly increases computation. Mixture-of-Experts models (MoEs) address this by allowing model capacity to scale without substantially increasing training or inference costs. Despite their promising results, MoE models encounter several challenges. Primarily, the dispersion of training tokens across multiple experts can lead to underfitting, particularly for infrequent tokens. Additionally, while fixed routing mechanisms can mitigate this issue, they compromise on the diversity of representations. In this paper, we propose MaskMoE, a method designed to enhance token-level learning by employing a routing masking technique within the Mixture-of-Experts model. MaskMoE is capable of maintaining representation diversity while achieving more comprehensive training. Experimental results demonstrate that our method outperforms previous dominant Mixture-of-Experts models in both perplexity (PPL) and downstream tasks.
1 Introduction
Large language models have achieved promising performance on various downstream natural language tasks Touvron et al. (2023a); Dai et al. (2022); Brown et al. (2020); Anil et al. (2023). However, training large language models on extensive textual data has significantly increased computational costs compared to previous works (e.g., BERT Devlin et al. (2019), ELMO Peters et al. (2018), and LSTM Hochreiter and Schmidhuber (1997)). In order to further scale up models within computational budgets, sparse activation networks Child et al. (2019); Du et al. (2022a) receive widespread attention due to their ability to significantly reduce computational costs by using only a few parameters per input. A widely studied approach is Mixture-of-Experts (MoEs)  Lepikhin et al. (2021); Du et al. (2022a); Dai et al. (2024); Fedus et al. (2022); Roller et al. (2021), which trains multiple expert networks but selects only a subset to process specific inputs. Compared to dense networks of the same model size, MoEs effectively reduce computational costs.
The current commonly used dynamic routing method in MoEs is to select the expert with the high confidence based on the probability distribution output by an intermediate layer with learnable parameters, i.e., a softmax layer. Previous works Lepikhin et al. (2021); Du et al. (2022a); Dai et al. (2024); Fedus et al. (2022) show that MoE models trained based on dynamic routing achieve better performance than dense models with the same amount of training computation.
However, this dynamic routing method introduces a new problem: the routing fluctuation problem Dai et al. (2022). This implies that as training progresses, the same tokens are assigned to different experts in different iterations. Consequently, each expert is trained on only a subset of the same tokens. Additionally, the lack of knowledge sharing and interaction among experts may negatively impact the model’s performance Zhao et al. (2024). Specifically, for the same tokens assigned to different experts, the knowledge each expert learns is harder to share. As previous work Shazeer et al. (2017) has indicated, when there are too many experts, the MoE model exhibits underfitting, and even shows higher perplexity than scenarios with fewer experts. The MoEs models have relatively less impact on the learning of frequent tokens, as these tokens have sufficient data to ensure that each expert receives adequate training. However, the MoEs model disperses infrequent tokens and concepts across various experts, which can lead to underfitting for these tokens. This issue hinders scaling the number of experts for the MoE model. As the number of experts increases while keeping the training data constant, the average number of tokens assigned to each expert further decreases, exacerbating the problem of expert underfitting.
Figure 1: Explanation of random routing masking strategy. For each token, some experts will be randomly masked, with only a subset of experts being visible, e.g., for the token “in” only experts 1, 2 and 3 are visible.
Previously fixed routing strategies Roller et al. (2021) based on Hash Layer, pre-assigning each token to a fixed-route expert, e.g., “deliberate” can only be routed to expert number one as shown in Figure 1. The fixed routing to send the same token to the same expert is more advantageous for the token’s thorough learning, especially for some infrequent tokens. However, previous works He et al. (2023); Yang et al. (2019) show that increasing the number of optional experts for the same token input enhances the diversity of representations. Therefore, fixed routing based on the Hash Layer may represent a deficiency in diversity, especially for the representation of frequent tokens. Frequent tokens, in order to achieve diverse representations, require encoding by more experts due to their broader range of usage Koranda et al. (2018). For instance, “the” can follow a variety of nouns, and having a more diverse representation helps to distinguish these nouns, making it easier to predict them in the next token prediction task. To enhance the diversity of representation for frequent tokens, it is necessary to employ more experts to encode them.
In this paper, we propose a routing strategy called MaskMoE. Specifically, we propose a routing masking method to adjust the number of visible experts for tokens of different frequencies by generating a masking vector for each token in the vocabulary before training. For infrequent tokens, the MoE layers employ routing masking to retain only one visible expert to which the token can be routed. For frequent tokens, the MoE layers have a lower routing masking rate, allowing more visible experts. MaskMoE can enhance token-level learning through a routing mask. Specifically, MaskMoE allows the model to learn more intensively about infrequent tokens by routing the same infrequent tokens to the same expert, enabling the expert to be fully trained on those tokens. Meanwhile, MaskMoE maintains the diversity of representations for frequent tokens by still having multiple experts available for routing, which benefits their representation diversity He et al. (2023); Yang et al. (2019). Additionally, despite frequent tokens being routed to multiple experts, their sufficient quantities ensure thorough training. Increase the visible number of experts for high-frequency tokens, i.e., increasing the count of updatable parameters, can further enhance capability. Our experimental results indicate that MaskMoE outperforms previous MoEs models, whether in terms of perplexity or downstream benchmark datasets.
Our contributions are summarized as follows:
  • We propose MaskMoE, which introduces a routing mask method to assign a different number of visible experts to tokens based on their frequency to enhance the model’s token-level learning. MaskMoE ensures sufficient training for infrequent tokens while maintaining diverse representations for frequent tokens.
  • We conduct extensive experiments to analyze the relationship between the number of experts and token frequency. We highlight that dynamic routing, which disperses the same tokens, can lead to underfitting of experts on infrequent tokens, while fixed routing lacks representation diversity for frequent tokens.
  • We validate the effectiveness of the proposed MaskMoE with extensive experiments. Experimental results show that it consistently outperforms previous Mixture-of-Experts models.
2 Related Work
2.1 Language Models
Language models are statistical models designed to optimize the likelihood of token sequences in training data Touvron et al. (2023a). Initially, language models relied on $n$-gram statistics Bahl et al. (1983); Katz (1987); Kneser and Ney (1995). Subsequently, the emphasis shifted to neural network-based models, particularly Recurrent Neural Networks Mikolov et al. (2010) and their variants like LSTMs Graves (2013). These models have demonstrated the ability to learn intricate patterns within textual data, achieving significant success in various language modeling applications.
In recent times, Transformers have become the predominant architecture for language models. Notable examples include BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), GPT-2 Radford et al. (2019), UniLM Dong et al. (2019), and T5 Raffel et al. (2020). The introduction of GPT-3 Brown et al. (2020), which boasts 175 billion parameters, marked a significant milestone due to its exceptional performance across numerous downstream tasks. This has led to a surge in research focusing on large generative language models, including prominent works such as Gopher Rae et al. (2021), PaLM Chowdhery et al. (2022), Pythia Biderman et al. (2023), OPT Zhang et al. (2022), GLM Du et al. (2022b); Zeng et al. (2023) and LLaMA Touvron et al. (2023a, b). Currently, GPT4 OpenAI (2023), trained with Reinforcement Learning from Human Feedback (RLHF), achieves truly remarkable results.
However, as the size of the model grows, the computational demands for both training and inference also increase. MoE models achieve scalability by sparsely activating portions of the model’s parameters, allowing for an increase in model size without significantly raising computational overhead. Consequently, MoE models have been receiving increasing attention recently.
2.2 Mixture-of-Experts
The concept of Mixture-of-Experts (MoE) models is initially proposed by  Jacobs et al. (1991). Later,  Shazeer et al. (2017) applies MoE to LSTM, training an LSTM model with up to 137B parameters. With the rise of the Transformer architecture Vaswani et al. (2017); Devlin et al. (2019), Gshard Lepikhin et al. (2021) first applies MoE to Transformers. Subsequently, powerful MoE models such as GLaM Du et al. (2022a) and Switch Transformer Fedus et al. (2022) emerge.
Early works  Zoph et al. (2022); Fedus et al. (2022); Du et al. (2022a); Lepikhin et al. (2021) mainly focus on dynamic routing methods for learning-to-route MoEs. The Hash Layer  Roller et al. (2021) proposes using a random hashing method to route tokens to a fixed expert, achieving better results than dynamic routing. Then,  Dai et al. 2022 points out that dynamic routing has the problem of routing fluctuation problem, meaning the same input is assigned to different experts as training progresses. Routing fluctuation often harms sample efficiency because the same input updates different experts, which significantly affects infrequent tokens. However, previous works He et al. (2023); Yang et al. (2019) indicate that having more experts allows tokens to obtain richer representations, and fixed routing can affect the diversity of representations, especially for frequent tokens.
Based on the above analysis, we propose MaskMoE, which uses routing masks to alleviate the underfitting problem of infrequent tokens caused by routing fluctuations, while maintaining the representational diversity of frequent tokens.
3 Method
In this section, we first review the training of the MoEs models. Then, we introduce detailed MaskMoE. Finally, we introduce the load balance loss.
3.1 Reviewing the Mixture-of-Experts
The routing of MoEs can be divided into fixed routing Roller et al. (2021); Dai et al. (2022) and dynamic routing with learnable parameters Lepikhin et al. (2021); Du et al. (2022a); Dai et al. (2024); Fedus et al. (2022). The common method for fixed routing is to use random hashing to determine Roller et al. (2021), before training, which experts each token will be sent to. In contrast, dynamic routing employs a routing layer with learnable parameters, i.e., a softmax layer, to decide which expert should process the input tokens.
In terms of structure, the FFN layer of the MoEs model is replaced with sparse expert modules. Given a tokenized input sequence $\mathbf{x}={x_{1},x_{2},...,x_{T}}$ consisting of $T$ tokens, the representation of each token is computed as hidden states by a standard Transformer.
$\mathbf{H}=Tansformer(Embedding(\mathbf{x}))$ (1)
Here, $\mathbf{H}$ denote the hidden states, i.e., $\mathbf{H}=[\mathbf{h}_{1},\mathbf{h}_{2},\ldots,\mathbf{h}_{T}]$. The dense transformer layer consists of self-attention and FFN sublayers.
$\mathbf{\bar{h}}_{t}^{l}=SelfAttn(\mathbf{h_{t}^{l-1}})$ (2)
$\mathbf{h_{t}^{l}}=FFN(\mathbf{\bar{h}}_{t}^{l})$ (3)
Here, $l$ represents the $l$-th transformer layer. MoEs methods replace the FFN sub-layer with the MoEs module. Each MoEs module consists of multiple FFNs, whose outputs are mixed with the gating function $P(\cdot)$, represented by the following formula.
$\mathbf{h_{t}^{l}}=\sum_{i}^{N}P(\mathbf{\bar{h}}_{t}^{l})\cdot FFN_{i}(\mathbf{\bar{h}}_{t}^{l})$ (4)
$N$ represents the number of experts in a single MoE layer. The majority of elements in the distribution of the gated $P(\cdot)$ output are zeros, i.e., only a small portion of experts will be activated. Therefore, increasing the total number of experts in MoEs does not increase computation or inference time.
3.2 Proposed MaskMoE
The typical dynamic routing method routes tokens to the top experts based on the scores output by a softmax layer. This approach to update expert parameters is sparse, meaning that the parameters of unselected experts are not updated. This sparsity may lead to experts underfitting for infrequent tokens. Fixed routing ensures that specific experts consistently process the same input tokens, which helps mitigate the problem of insufficient learning for infrequent tokens. However, it is less effective in fostering a diversity of representations compared to dynamic routing methods. Dynamic routing allows the same token to be encoded by different experts, leading to more varied representations He et al. (2023); Yang et al. (2019). Since frequent tokens can have different meanings depending on the context, fixed routing is less effective than dynamic routing in capturing the nuances of these tokens. To balance the density of parameter updates and the diversity of representations in MoE models, we propose MaskMoE to boost token-level learning via routing mask. As illustrated in Figure 1, MaskMoE achieves this by introducing a routing masking mechanism. For each token, a fixed random masking is applied to the experts, allowing them to be routed only to their respective subsets of experts. The formula is represented as follows.
$\mathbf{p}$ $=\text{Softmax}(W_{g}\mathbf{\bar{h}_{t}^{l}}+\mathbf{M})$ (5)
Where $\mathbf{M}$ is the masking vector used to control the visibility of token-specific experts. For visible experts, the corresponding elements in $\mathbf{M}$ are 0, while for invisible experts, the corresponding elements in $\mathbf{M}$ are $\infty$- ∞. The masking vector is determined before training and does not change during the training process. For multi-layer MoE models, we reuse the same masking vector $\mathbf{M}$ across different MoE layers. Here, $\mathbf{W_{g}}$ is the parameter of the gate. The initial energy masking vector for a token $t$ can be represented by the following formula.
$\mathbf{M}^{t}$ $:=-\infty\cdot\mathbf{1}_{N}$ (6)
$\mathbf{C}$ =$=\{\text{C}_{i}\}_{i=1}^{V}\sim\mathcal{U}(\{0:N-1\})$
$\mathbf{M}_{j}^{t}$ $=0,\quad\forall j\in\mathbf{C}$
Formally, we denote the number of visible experts as $V$. $\mathcal{U}(\{0:N-1\})$ denotes a uniform distribution over the set of integers from 00 to $N-1$. It is worth noting that for infrequent tokens, each token has only one visible expert, i.e., $V=1$. This promotes more thorough learning by the selected expert for such infrequent tokens. For frequent tokens, we also reduce the number of visible experts, but the number of visible experts is greater than 1, i.e., $1<V<=N$. For frequent tokens, it is common for $V$ to be less than $N$. We believe that an appropriately sized $V$ not only allows frequent words to have higher representation diversity compared to fixed routing but also enables more thorough training for the same token compared to dynamic routing.
3.3 Load Balance Loss
MoE models typically require distributed training, where different experts are deployed across various nodes. To avoid load imbalance, where a minority of expert nodes handle the majority of tokens while the majority of expert nodes spend more time idle, this can impact training efficiency. It is generally desirable for the number of tokens processed by different experts to be roughly equal. To achieve load balancing, a load balancing loss is commonly introduced in the training of MoE models. We follow previous works Huang et al. (2024); Fedus et al. (2022) and adopt a widely used loss.
$\mathcal{L}_{bal}$ $=N*\sum_{i=1}^{N}w_{i}\cdot R_{i}$ (7)
[…]
$R_{i}$ $=\frac{1}{T}\sum_{j=1}^{T}\mathbf{p}_{i}^{j}$ (9)
Where $T$ represents the number of tokens in a mini-batch. $\mathbf{p}^{j}$ denotes the probability distribution of the routing output for the j-th token, while $\mathbf{p}_{i}^{j}$ represents the probability value of the i-th expert. It is noteworthy that for infrequent tokens, which have only one visible expert, the load balancing loss does not apply to them. The balancing loss primarily regulates the routing of frequent tokens. Additionally, since infrequent tokens are assigned to experts through random hashing, the overall load is relatively balanced. Therefore, our experiments find that MaskMoE, compared to SMoE, does not experience a significant decrease in training efficiency.
Our final loss is a combination of the language model loss and load-balance loss:
$\mathcal{L}=\mathcal{L}_{lm}+\mathcal{L}_{bal}$ (10)
4 Expertiments
Model Configuration Params Activated Params Pile PPL
Standard Transformer Layers=24, Dense 468M 468M 6.95
SMoE Layers=24, MoE_Layers = 1 1.3B 468M 6.62
Hash Layer Layers=24, MoE_Layers = 1 1.3B 468M 6.56
Share-MoE Layers=24, MoE_Layers = 1 1.3B 468M 6.72
MaskMoE Layers=24, MoE_Layers = 1 1.3B 468M 6.48
SMoE Layers=24, MoE_Layers = 12 10B 468M 6.18
Hash Layer Layers=24, MoE_Layers = 12 10B 468M 6.16
Share-MoE Layers=24, MoE_Layers = 12 10B 468M 6.15
MaskMoE Layers=24, MoE_Layers = 12 10B 468M 6.11
Table 1: Perplexity results of language modeling.
4.1 Pre-training Dataset
Following previous works, we use the Pile dataset Gao et al. (2021) for pre-training data. The Pile dataset is a large-scale, publicly available corpus for language model pre-training, containing 22 domains and over 825GB of English text data. For our experiments, we use the well-known LLaMA tokenizer for tokenization, with a vocabulary size of $32k$. We follow  Xie et al. (2023a); Su et al. (2024) to calculate the sampling rate for each domain based on the number of tokens after tokenization. Due to the computational budget, we follow the pretraining settings of  Xie et al. (2023a); Su et al. (2024); Huang et al. (2024); Xiong et al. (2024); Lian et al. (2024), all models are pre-trained with 100B tokens.
Then, to identify infrequent tokens and frequent tokens, we calculate the frequency of each token in the training set and sort them in descending order of frequency. We categorize the top tokens that cover $P$of the dataset as frequent tokens, and the remaining $(1-P)$as infrequent tokens. Where $P$ is a tunable hyperparameter.
4.2 Experimental Setup
Following  Touvron et al. (2023a, b); Huang et al. (2024), we adopt the LLaMA architecture with 24 transformer layers and a hidden layer dimension of 1024. In line with LLaMA Touvron et al. (2023a), we employ the AdamW  Loshchilov and Hutter (2019) optimizer, and applying a cosine learning rate decay schedule.
For the MoE model, following  Roller et al. (2021), we conduct experiments under both single-layer and multi-layer settings. For the single-layer setting, we replace the final FFN layer with the MoE layer. For the multi-layer MoE models, we follow Gshard Lepikhin et al. (2021), alternating MoE layers every other layer, resulting in a total of $12$ MoE layers in this setting. For single-layer MoE models and dense models, we use a learning rate of $43e-$ 4, while for $12$-layer MoE models, we employ a learning rate of $41e-$ 4 to ensure stable convergence of the models. If not specified otherwise, tokens that account for the top 40% of frequency are considered frequent tokens, while the remaining are considered infrequent tokens, i.e., $P=0.4$. For frequent tokens, the visible expert count $V$ is set to 8, and for infrequent tokens, the visible expert count $V$ is set to 1. All of our implementations are based on the DeepSpeed [^1] https://github.com/microsoft/DeepSpeed library Rajbhandari et al. (2022); Rasley et al. (2020), which offers robust support for MoE-distributed training. By default, we enable the random token selection strategy Kim et al. (2021) implemented in the library to facilitate faster model convergence and better runtime efficiency. Following previous works Roller et al. (2021); Fedus et al. (2022); Dai et al. (2022), unless otherwise specified, our experiments only select the top expert.
4.3 Compared Models
Including MaskMoE, we compare five models for validation experiments. Next, we will briefly introduce five compared models.
  • Dense represents a standard Transformer language model with 468M parameters.
  • SMoE denotes a Switch Transformer Fedus et al. (2022), where the router is a learnable parameter. Unless otherwise specified, each MoE layer has 64 experts.
  • Hash Layer Roller et al. (2021) signifies that through a random hash method, each token is assigned a fixed routing expert prior to training. Unless otherwise specified, each MoE layer has 64 experts.
  • Share-MoE is a hybrid dense and MoE model created using residual connections Rajbhandari et al. (2022). Models with shared experts are currently a popular architecture Dai et al. (2024); DeepSeek-AI et al. (2024); Zhao et al. (2024); Rajbhandari et al. (2022). Unless otherwise specified, Share-MoE has 1 shared expert and 128 routed experts, where each expert is 0.5 times the size of a standard FFN. During both training and inference, in addition to the activation of shared experts, the top expert is also selected from a pool of 128 experts. In such a setup, Share-MoE has the same number of floating-point operations (FLOPs) and activation parameters as SMoE and Hash Layer. This allows for a fair comparison between Share-MoE, SMoE, and Hash Layer.
  • MaskMoE achieves a balance between representation diversity in MoE layers and thorough parameter training through the routing masking method we proposed. Similarly, in this section, MaskMoE also employs the shared expert architecture, which is identical to that of Share-MoE.
Model BoolQ Hellaswag LAMBADA PIQA SIQA StoryCloze Arc-e TriviaQA WebQs
Standard Transformer 56.02 40.73 52.55 67.62 40.79 63.55 51.43 7.44 4.97
MoE_Layers = 1
SMoE 55.26 43.66 53.46 68.28 41.15 64.14 51.98 10.00 6.99
Hash Layer 56.61 43.88 54.67 69.53 41.76 64.19 53.70 9.82 6.89
Share-MoE 58.87 43.00 53.00 68.39 41.45 64.19 51.92 10.54 6.89
MaskMoE 58.38 44.47 55.36 68.99 41.86 65.15 53.14 10.39 7.14
MoE_Layers = 12
SMoE 56.18 46.92 56.37 69.91 41.20 64.99 54.34 13.54 7.33
Hash Layer 57.16 46.61 56.68 71.00 40.83 65.47 55.05 12.97 5.95
Share-MoE 58.71 46.94 56.39 69.69 40.84 65.42 55.89 12.77 6.94
MaskMoE 58.32 47.46 57.46 70.62 41.91 65.69 55.35 15.20 6.74
Table 2: Performances of language models on downstream benchmarks. The best score is marked in bold, and the second best is underlined.
4.4 Main Results
In this section, we first present the model’s perplexity (PPL) on the Pile validation set. Then, following Touvron et al. (2023a); Brown et al. (2020); Su et al. (2024); Dai et al. (2024), we conduct tests on a large number of downstream benchmarks, including zero-shot for BoolQ Clark et al. (2019), HellaSwag Zellers et al. (2019), LAMBADA Paperno et al. (2016), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), StoryCloze Mostafazadeh et al. (2016), and Arc-e Bhakthavatsalam et al. (2021), as well as 5-shot for TriviaQA Joshi et al. (2017) and WebQs Berant et al. (2013). Among them, TriviaQA and WebQs use exact match as the metric, while the remaining benchmarks are evaluated based on accuracy. For a fair comparison, we use the open-source evaluation tool lm-evaluation-harness [^2] https://github.com/EleutherAI/lm-evaluation-harness for assessment.
4.4.1 Perplexity Results
Table 1 shows the main results of language modeling on the Pile validation sets. With the same parameter activations maintained, MoE models consistently achieve improvements compared to the dense model. In comparison to SMoE and Hash Layer, under a single MoE layer setup, MaskMoE has reduced the PPL on the Pile Validation by 0.14 and 0.08 respectively. In a 12-MoE layers setup, the PPL reductions are 0.07 and 0.05 respectively. Compared to Share-MoE, MaskMoE decreases by 0.24 in single MoE layers and by 0.04 in 12-MoE layers. These results demonstrate the rationality of the MaskMoE design.
It is noteworthy that Share-MoE achieves worse results than SMoE and Hash Layer in the single MoE setup. However, MaskMoE consistently maintains superior performance, whether in single or multi-MoE layers. This indicates that MaskMoE exhibits stronger robustness across various settings.
4.4.2 Benchmark Results
As shown in Table 2, we present the performance of the models on downstream tasks. Firstly, we observe that MoE models significantly outperform Dense models in downstream tasks. Then, we find that MaskMoE substantially outperforms the other MoE baselines, whether in single-layer or 12-layers settings.
Compared to SMoE, MaskMoE outperforms in all 9 benchmarks with a single MoE layer and in 8 out of 9 benchmarks with 12-MoE layers. Against Hash Layer, MaskMoE excels in 7 out of 9 benchmarks with a single MoE layer and in 8 out of 9 with 12-MoE layers. Compared to Share-MoE, MaskMoE shows significant improvements in 7 out of 9 benchmarks with a single MoE layer and in 6 out of 9 with 12-MoE layers.
This indicates that the performance enhancement is achieved by using our proposed router mask to control the visibility of experts for different tokens. We believe the improvements are due to two factors. Firstly, compared to SMoE and Share-MoE, infrequent tokens are only routed to a single expert, ensuring more thorough training for these tokens. Secondly, in contrast to the Hash Layer, frequent tokens have multiple expert options, which promotes diversity in the representation of frequent tokens.
5 Analyses
We conduct further experiments to provide more insightful analyses on the proposed MaskMoE. Considering the constraints of computational resources, following previous works Roller et al. (2021); Xie et al. (2023b); Dai et al. (2022), unless otherwise specified, most of the analytical experiments are conducted under the setting of a single MoE layer.
5.1 Impact of Router Mask
$P$ $V_{a}$ $V_{b}$ PPL $P$ $V_{a}$ $V_{b}$ PPL
0.2 8 1 6.518 0.4 16 1 6.509
0.4 8 1 6.506 0.4 4 1 6.511
0.6 8 1 6.518 0.4 64 1 6.549
/ 64 64 6.618$\spadesuit$ / 1 1 6.558 $\clubsuit$
/ 4 4 6.551 / 8 8 6.566
Table 3: Impact of Router Mask. $\spadesuit$ denotes SMoE, $\clubsuit$ denotes Hash Layer. To make the differences in MaskMoE under various configurations more apparent, the values shown in the table are precise to three decimal places.
To visually demonstrate the impact of the routing masking mechanism, the experiments in this subsection do not employ a shared expert structure. Instead, they adopt the same structure as SMoE and Hash Layer, with each MoE layer comprising 64 experts, and the FNN layer is full-sized.
The performance of MaskMoE is predominantly influenced by two hyperparameters: the boundary threshold $P$ for the division of frequent and infrequent tokens, and the maximum number of visible experts $V$ for tokens. As shown in Table 3, we conduct an experimental parameter search for both and reported the Pile Validation PPL. $V_{a}$ represents the number of experts visible for frequent tokens, and $V_{b}$ represents the number of experts visible for infrequent tokens. Notably, when $V_{a}=V_{b}=1$, MaskMoE is equivalent to the Hash Layer, and when $V_{a}=V_{b}=64$, MaskMoE is equivalent to SMoE.
We have the following observations. Firstly, we find that MaskMoE outperforms the Hash Layer. With the $V_{a}=8$ and $V_{b}=1$ settings, MaskMoE reduces perplexity by 0.052 on the Pile validation set compared to the Hash Layer, i.e., $V_{a}=V_{b}=1$. This indicates that MaskMoE’s use of multiple expert encodings for frequent tokens enhances representation diversity, which is beneficial for model performance.
Secondly, compared to SMoE, concentrating tokens into fewer visible experts, such as settings like $V_{a}=V_{b}=8$ or $V_{a}=V_{b}=4$, yields better outcomes. This indicates that routing tokens more densely to a reduced number of visible experts enhances thorough training and improves model performance. Moreover, we observe that the configuration $V_{a}=8$ and $V_{b}=1$ outperforms $V_{a}=V_{b}=8$. This indicates that further reducing the number of visible experts is more beneficial for infrequent tokens.
Lastly, it is noteworthy that for MaskMoE, when $V_{b}=1$, regardless of the value of $V_{a}$ (whether it is 4, 8, 16, or 64), the performance of MaskMoE consistently outperforms that of SMoE and Hash Layer. This demonstrates the robustness of MaskMoE.
5.2 Impact of Shared Expert
To explore the impact of shared experts, in this subsection, we conduct ablation experiments on shared experts, and the results are shown in Table 4. We separately report the average scores of the benchmarks and the PPL of Pile Validation. It is noteworthy that after removing the shared experts, we adopt the same architecture as SMoE and Hash Layer, where the MoE layer consists of 64 fully-sized FFN.
We find that even without shared experts, MaskMoE outperforms SMoE, Hash Layer, and Share-MoE in average benchmark scores and Pile Validation PPL. This further validates the soundness of MaskMoE’s design. Additionally, we observe that incorporating shared expert structures continuously improves MaskMoE’s performance. We believe this is because the shared expert structure promotes greater specialization in token representation DeepSeek-AI et al. (2024).
Model Avg($\uparrow$) PPL($\downarrow$)
SMoE 43.87 6.62
Hash Layer 44.56 6.56
Share-MoE 44.25 6.72
MaskMoE 44.99 6.48
w/o Shared Experts 44.85 6.51
Table 4: Average scores of benchmarks and perplexity on the Pile validation. The best score is marked in bold, and the second best is underlined.
5.3 Number of Experts
(a) Non-Shared Expert Structures
(b) Shared Expert Structures
Figure 2: Comparison of MoE-based Transformers with different numbers of experts.
To investigate the variation in model performance under different configurations of expert numbers, we conducted experiments with 16, 64, and 128 experts. The experimental results, as shown in Figure 2, reveal the following observations.
Firstly, we observe that for SMoE and Share-MoE, increasing the number of experts does not necessarily lead to a lower perplexity. In fact, when the number of experts reaches 128, the model exhibits a higher PPL compared to having only 64 experts. We believe this is because with more experts, the average number of tokens assigned to each expert decreases, leading to further underfitting of the model. In contrast, the performance of the Hash Layer and MaskMoE consistently improves with the increase in the number of experts, indicating that they are more favorable for scaling the number of experts.
Secondly, we observe that compared to the Hash Layer, MaskMoE consistently performs better across all configurations. Although the Hash Layer also enables experts to be adequately trained with the tokens assigned to them, MaskMoE excels in the diversity of representations for frequent tokens. This superiority in handling token diversity contributes to MaskMoE’s overall advantage over Hash Layer.
Lastly, we find that regardless of the number of experts, MaskMoE consistently maintains the best performance. This indicates that MaskMoE has strong parameter robustness.
6 Conclusions
In this paper, we analyze dynamic routing models such as SMoE, which routes the same token to multiple experts, potentially causing the model to underfit, especially for infrequent tokens. While fixed routing can mitigate this issue, it struggles with capturing rich features for frequent tokens. To address these issues, we propose a routing masking method aimed at ensuring adequate learning of infrequent tokens while maintaining rich representations for frequent tokens. Extensive experiments demonstrate that the proposed MaskMoE outperforms various downstream tasks and significantly improves perplexity (PPL) on the Pile dataset compared to SMoE, Hash Layer, and Share-MoE.
7 Limitiations
In this paper, we categorize tokens simply into frequent and infrequent tokens by their frequency. This classification seems somewhat rigid. Considering a smoother classification method could potentially yield better results. Specifically, when tokens are arranged in ascending order, their frequencies demonstrate a gradual decline, and the corresponding number of visible experts also decreases progressively. Due to computational constraints, we do not conduct further experiments. We leave the exploration of smoother partitioning methods for future work.

### v2

@@
 Abstract
-Scaling model capacity enhances its capabilities but significantly increases computation.
-Mixture-of-Experts models (MoEs) address this by allowing model capacity to scale without substantially increasing training or inference costs.
+Scaling the size of a model enhances its capabilities but significantly increases computation complexity.
+Mixture-of-Experts models (MoE) address the issue by allowing model size to scale up without substantially increasing training or inference costs.
 Despite their promising results, MoE models encounter several challenges.
-Primarily, the dispersion of training tokens across multiple experts can lead to underfitting, particularly for infrequent tokens.
-Additionally, while fixed routing mechanisms can mitigate this issue, they compromise on the diversity of representations.
+Primarily, for dynamic routing methods, the dispersion of training tokens across multiple experts can lead to underfitting, particularly for infrequent tokens.
+Additionally, while fixed routing methods can mitigate that issue, they compromise on the diversity of representations.
 In this paper, we propose MaskMoE, a method designed to enhance token-level learning by employing a routing masking technique within the Mixture-of-Experts model.
 MaskMoE is capable of maintaining representation diversity while achieving more comprehensive training.
-Experimental results demonstrate that our method outperforms previous dominant Mixture-of-Experts models in both perplexity (PPL) and downstream tasks.
+Experimental results demonstrate that our method outperforms previous dominant Mixture-of-Experts models in terms of both perplexity (PPL) and downstream task performance.
 1 Introduction
@@
 (2023).
-However, training large language models on extensive textual data has significantly increased computational costs compared to previous works (e.g., BERT Devlin et al.
-(2019), ELMO Peters et al.
-(2018), and LSTM Hochreiter and Schmidhuber (1997)).
+Moreover, according to the scaling law  Kaplan et al.
+(2020); Hoffmann et al.
+(2022), as the model size increases, the model’s size will continue to grow.
+However, for dense language models, the computational costs of continuing to scale up are substantial.
 In order to further scale up models within computational budgets, sparse activation networks Child et al.
 (2019); Du et al.
-(2022a) receive widespread attention due to their ability to significantly reduce computational costs by using only a few parameters per input.
-A widely studied approach is Mixture-of-Experts (MoEs)  Lepikhin et al.
+(2022a) receive widespread attention due to their ability to significantly reduce computational costs by using only a part of parameters per input.
+A widely studied approach is Mixture-of-Experts (MoE)  Lepikhin et al.
 (2021); Du et al.
@@
 (2021), which trains multiple expert networks but selects only a subset to process specific inputs.
-Compared to dense networks of the same model size, MoEs effectively reduce computational costs.
-The current commonly used dynamic routing method in MoEs is to select the expert with the high confidence based on the probability distribution output by an intermediate layer with learnable parameters, i.e., a softmax layer.
+Compared to dense networks of the same model size, MoE effectively reduces computational costs and achieves better results in both PPL and downstream tasks Lepikhin et al.
+(2021); Du et al.
+(2022a); Dai et al.
+(2024).
+In MoE models, for each input (e.g., a token in the language models), the router in front of the experts needs to decide which tokens to distribute to which experts.
+The current commonly used dynamic routing method in MoE is to select the expert with the top $k$ highest confidence based on a probability distribution output by an intermediate layer with learnable parameters.
 Previous works Lepikhin et al.
@@
 (2022) show that MoE models trained based on dynamic routing achieve better performance than dense models with the same amount of training computation.
-However, this dynamic routing method introduces a new problem: the routing fluctuation problem Dai et al.
+However, this dynamic routing method introduces a new challenge: the routing fluctuation problem Dai et al.
 (2022).
-This implies that as training progresses, the same tokens are assigned to different experts in different iterations.
-Consequently, each expert is trained on only a subset of the same tokens.
-Additionally, the lack of knowledge sharing and interaction among experts may negatively impact the model’s performance Zhao et al.
+It implies that as training moves on, the same token is assigned to different experts in different iterations.
+This has two negative impacts.
+Firstly, each expert is trained on only a subset of the occurrences of an identical token, resulting in each expert not learning enough about this token.
+As previous work Shazeer et al.
+(2017) has indicated, when there are too many experts, the MoE model exhibits underfitting, and even shows higher PPL than its counterparts with fewer experts.
+The routing fluctuation problem has relatively less impact on the learning of frequent tokens, as those tokens have sufficient occurrences to ensure that each expert receives adequate training.
+On the contrary, for infrequent tokens, the routing fluctuation problem can disperse each of them across various experts, which can lead to underfitting for them.
+Secondly, the knowledge each expert learns from the corresponding occurrence subset w.r.t an identical token is hard to share among experts.
+Such lack of knowledge sharing and interaction among experts may negatively impact the model’s performance Zhao et al.
 (2024).
-Specifically, for the same tokens assigned to different experts, the knowledge each expert learns is harder to share.
-As previous work Shazeer et al.
-(2017) has indicated, when there are too many experts, the MoE model exhibits underfitting, and even shows higher perplexity than scenarios with fewer experts.
-The MoEs models have relatively less impact on the learning of frequent tokens, as these tokens have sufficient data to ensure that each expert receives adequate training.
-However, the MoEs model disperses infrequent tokens and concepts across various experts, which can lead to underfitting for these tokens.
-This issue hinders scaling the number of experts for the MoE model.
-As the number of experts increases while keeping the training data constant, the average number of tokens assigned to each expert further decreases, exacerbating the problem of expert underfitting.
-Figure 1: Explanation of random routing masking strategy.
-For each token, some experts will be randomly masked, with only a subset of experts being visible, e.g., for the token “in” only experts 1, 2 and 3 are visible.
-Previously fixed routing strategies Roller et al.
-(2021) based on Hash Layer, pre-assigning each token to a fixed-route expert, e.g., “deliberate” can only be routed to expert number one as shown in Figure 1.
-The fixed routing to send the same token to the same expert is more advantageous for the token’s thorough learning, especially for some infrequent tokens.
-However, previous works He et al.
+The above two issues hinder scaling the number of experts for an MoE model.
+Because, as the number of experts increases while keeping the training data constant, the average number of tokens assigned to each expert further decreases, exacerbating the above two issues.
+Figure 1: Illustration of random routing masking method.
+For each token, some experts will be randomly masked, with only a subset of experts being visible, e.g., for the token “in” only expert FFN #1, #2, and #3 are visible.
+To alleviate the issue of routing fluctuations, the previously proposed fixed routing method Roller et al.
+(2021); Dai et al.
+(2022) is a solution.
+Specifically, Hash Layer Roller et al.
+(2021) pre-assigns the same token to a fixed routing expert.
+For instance, as shown in Figure 1, the token “deliberate” can only be routed to the expert FFN#2.
+The fixed routing method to send all the occurrences of a token to the same expert is advantageous for the token’s thorough learning, especially for infrequent tokens.
+However, fixed routing based on the Hash Layer represents a deficiency in diversity, especially for those frequent tokens.
+Previous works He et al.
 (2023); Yang et al.
-(2019) show that increasing the number of optional experts for the same token input enhances the diversity of representations.
-Therefore, fixed routing based on the Hash Layer may represent a deficiency in diversity, especially for the representation of frequent tokens.
-Frequent tokens, in order to achieve diverse representations, require encoding by more experts due to their broader range of usage Koranda et al.
+(2019) show that increasing the number of optional experts for all occurrences of an identical can enhance the diversity of representations.
+Frequent tokens probably require to be encoded by more experts due to their broader range of usage Koranda et al.
 (2018).
-For instance, “the” can follow a variety of nouns, and having a more diverse representation helps to distinguish these nouns, making it easier to predict them in the next token prediction task.
-To enhance the diversity of representation for frequent tokens, it is necessary to employ more experts to encode them.
-In this paper, we propose a routing strategy called MaskMoE.
-Specifically, we propose a routing masking method to adjust the number of visible experts for tokens of different frequencies by generating a masking vector for each token in the vocabulary before training.
-For infrequent tokens, the MoE layers employ routing masking to retain only one visible expert to which the token can be routed.
-For frequent tokens, the MoE layers have a lower routing masking rate, allowing more visible experts.
-MaskMoE can enhance token-level learning through a routing mask.
-Specifically, MaskMoE allows the model to learn more intensively about infrequent tokens by routing the same infrequent tokens to the same expert, enabling the expert to be fully trained on those tokens.
+For instance, the token “the” appears in various contexts and owns different semantic meanings, and having a more diverse representation helps to distinguish it.
+Therefore, based on the observations above, we need a fixed routing strategy for infrequent tokens, while also needing more experts for frequent ones.
+To meet the requirements above.
+In this paper, we propose a routing method called MaskMoE, which includes a routing masking method that adjusts the number of visible experts for tokens of different frequencies, by generating a masking vector for each token in the vocabulary before training.
+Specifically, for each infrequent token, the proposed MaskMoE employs routing masking to retain only one visible expert to which the token can be routed, in each MoE layer.
+For each frequent token, the proposed MaskMoE has a lower routing masking rate to allow more visible experts for it in each MoE layer.
+MaskMoE can enhance token-level learning through such a routing mask.
+Specifically, MaskMoE makes the model learn more intensively about infrequent tokens by routing each to an identical expert in each MoE layer, enabling the expert to be fully trained w.r.t that token.
 Meanwhile, MaskMoE maintains the diversity of representations for frequent tokens by still having multiple experts available for routing, which benefits their representation diversity He et al.
@@
 (2019).
-Additionally, despite frequent tokens being routed to multiple experts, their sufficient quantities ensure thorough training.
-Increase the visible number of experts for high-frequency tokens, i.e., increasing the count of updatable parameters, can further enhance capability.
-Our experimental results indicate that MaskMoE outperforms previous MoEs models, whether in terms of perplexity or downstream benchmark datasets.
+Despite frequent tokens being routed to multiple experts, their sufficient quantities can still ensure thorough training.
+Our experimental results indicate that MaskMoE outperforms previous MoE models, in terms of either PPL or downstream task performance.
 Our contributions are summarized as follows:
-  • We propose MaskMoE, which introduces a routing mask method to assign a different number of visible experts to tokens based on their frequency to enhance the model’s token-level learning.
+  • We propose MaskMoE, which introduces a routing masking method to assign different numbers of visible experts to tokens based on their frequency to enhance token-level learning.
 MaskMoE ensures sufficient training for infrequent tokens while maintaining diverse representations for frequent tokens.
-  • We conduct extensive experiments to analyze the relationship between the number of experts and token frequency.
-We highlight that dynamic routing, which disperses the same tokens, can lead to underfitting of experts on infrequent tokens, while fixed routing lacks representation diversity for frequent tokens.
+  • We highlight that dynamic routing, which disperses the occurrences of a token into different experts during training, can lead to the underfitting of experts on infrequent tokens, while fixed routing lacks representation diversity for frequent tokens.
   • We validate the effectiveness of the proposed MaskMoE with extensive experiments. Experimental results show that it consistently outperforms previous Mixture-of-Experts models.
@@
 (2010) and their variants like LSTMs Graves (2013).
-These models have demonstrated the ability to learn intricate patterns within textual data, achieving significant success in various language modeling applications.
+Those models have demonstrated the ability to learn intricate patterns within textual data, achieving significant success in various language modeling applications.
 In recent times, Transformers have become the predominant architecture for language models.
@@
 (2020), which boasts 175 billion parameters, marked a significant milestone due to its exceptional performance across numerous downstream tasks.
-This has led to a surge in research focusing on large generative language models, including prominent works such as Gopher Rae et al.
+And that then led to a surge in research focusing on large generative language models, including prominent works such as Gopher Rae et al.
 (2021), PaLM Chowdhery et al.
@@
 However, as the size of the model grows, the computational demands for both training and inference also increase.
-MoE models achieve scalability by sparsely activating portions of the model’s parameters, allowing for an increase in model size without significantly raising computational overhead.
+MoE models achieve scalability by sparsely activating a portion of the model’s parameters, allowing for an increase in model size without significantly raising computational overhead.
 Consequently, MoE models have been receiving increasing attention recently.
 2.2 Mixture-of-Experts
-The concept of Mixture-of-Experts (MoE) models is initially proposed by  Jacobs et al.
+The concept of Mixture-of-Experts (MoE) models was initially proposed by  Jacobs et al.
 (1991).
 Later,  Shazeer et al.
-(2017) applies MoE to LSTM, training an LSTM model with up to 137B parameters.
+(2017) applied MoE to LSTM, training an LSTM model with up to 137B parameters.
 With the rise of the Transformer architecture Vaswani et al.
@@
 (2019), Gshard Lepikhin et al.
-(2021) first applies MoE to Transformers.
+(2021) further applied MoE to Transformers.
 Subsequently, powerful MoE models such as GLaM Du et al.
 (2022a) and Switch Transformer Fedus et al.
-(2022) emerge.
+(2022) emerged.
 Early works  Zoph et al.
@@
 (2022a); Lepikhin et al.
-(2021) mainly focus on dynamic routing methods for learning-to-route MoEs.
-The Hash Layer  Roller et al.
-(2021) proposes using a random hashing method to route tokens to a fixed expert, achieving better results than dynamic routing.
-Then,  Dai et al.
-2022 points out that dynamic routing has the problem of routing fluctuation problem, meaning the same input is assigned to different experts as training progresses.
-Routing fluctuation often harms sample efficiency because the same input updates different experts, which significantly affects infrequent tokens.
-However, previous works He et al.
+(2021) mainly focused on dynamic routing methods for learning-to-route MoE.
+Then the Hash Layer  Roller et al.
+(2021) is proposed to use a random hashing method to route tokens to a fixed expert, which even achieved better results than dynamic routing.
+As an explanation to the advantage of the simple fixed routing method,  Dai et al.
+2022 pointed out that dynamic routing has the problem of routing fluctuation problem, meaning that an identical input is assigned to different experts as training progresses.
+Routing fluctuation often harms sample efficiency, especially for the learning of infrequent tokens in the language model.
+On the contrary, other previous works He et al.
 (2023); Yang et al.
-(2019) indicate that having more experts allows tokens to obtain richer representations, and fixed routing can affect the diversity of representations, especially for frequent tokens.
-Based on the above analysis, we propose MaskMoE, which uses routing masks to alleviate the underfitting problem of infrequent tokens caused by routing fluctuations, while maintaining the representational diversity of frequent tokens.
+(2019) also argued that having more experts allows tokens to obtain richer representations, and fixed routing can affect the diversity of representations, especially for frequent tokens.
+Considering both routing fluctuation and representation diversity, we propose MaskMoE, which uses routing masks to alleviate the underfitting problem of infrequent tokens caused by routing fluctuations, while maintaining the representational diversity of frequent tokens.
 3 Method
-In this section, we first review the training of the MoEs models. Then, we introduce detailed MaskMoE. Finally, we introduce the load balance loss.
+In this section, we first review the training of the MoE models. Then, we elaborate on the proposed MaskMoE. Finally, we introduce the load balance loss.
 3.1 Reviewing the Mixture-of-Experts
-The routing of MoEs can be divided into fixed routing Roller et al.
+The routing of MoE can be divided into fixed routing Roller et al.
 (2021); Dai et al.
@@
 (2021), before training, which experts each token will be sent to.
-In contrast, dynamic routing employs a routing layer with learnable parameters, i.e., a softmax layer, to decide which expert should process the input tokens.
-In terms of structure, the FFN layer of the MoEs model is replaced with sparse expert modules.
-Given a tokenized input sequence $\mathbf{x}={x_{1},x_{2},...,x_{T}}$ consisting of $T$ tokens, the representation of each token is computed as hidden states by a standard Transformer.
-$\mathbf{H}=Tansformer(Embedding(\mathbf{x}))$ (1)
-Here, $\mathbf{H}$ denote the hidden states, i.e., $\mathbf{H}=[\mathbf{h}_{1},\mathbf{h}_{2},\ldots,\mathbf{h}_{T}]$. The dense transformer layer consists of self-attention and FFN sublayers.
-$\mathbf{\bar{h}}_{t}^{l}=SelfAttn(\mathbf{h_{t}^{l-1}})$ (2)
-$\mathbf{h_{t}^{l}}=FFN(\mathbf{\bar{h}}_{t}^{l})$ (3)
-Here, $l$ represents the $l$-th transformer layer.
-MoEs methods replace the FFN sub-layer with the MoEs module.
-Each MoEs module consists of multiple FFNs, whose outputs are mixed with the gating function $P(\cdot)$, represented by the following formula.
-$\mathbf{h_{t}^{l}}=\sum_{i}^{N}P(\mathbf{\bar{h}}_{t}^{l})\cdot FFN_{i}(\mathbf{\bar{h}}_{t}^{l})$ (4)
-$N$ represents the number of experts in a single MoE layer.
-The majority of elements in the distribution of the gated $P(\cdot)$ output are zeros, i.e., only a small portion of experts will be activated.
-Therefore, increasing the total number of experts in MoEs does not increase computation or inference time.
+In contrast, dynamic routing employs a routing layer with learnable parameters to decide which expert should process an input token.
+In terms of structure, the FFN layer of a dense language model is generally replaced with sparse expert modules to build an MoE model.
+Given a tokenized input sequence $\mathbf{x}={x_{1},x_{2},...,x_{T}}$ consisting of $T$ tokens , and then generates a probability distribution $\mathbf{p}_{v}$ over the vocabulary as output.
+In nearly all implementations, the Cross-Entropy loss is used as the loss function to maximize the predicted probability $\mathbf{p}_{v}^{x_{i}}$ w.r.t the ground-truth token $x_{i}$.
+The training loss $\mathcal{L}_{lm}$ of the generative language model can be formulated as follows.
+$\mathcal{L}_{lm}$ $=-\sum_{i=2}^{T}\log(\mathbf{p}_{v}^{x_{i}})$ (1)
+$s.t.,\mathbf{p}_{v}$ $=\text{softmax}(W\mathbf{H}^{L})$ (2)
+$\mathbf{H^{L}}$ $=\text{Transformer}(x_{1}^{0},x_{2}^{0},\ldots,x_{T}^{0})$ (3)
+Here, $\mathbf{H^{L}}=[\mathbf{h}_{1}^{L},\mathbf{h}_{2}^{L},\ldots,\mathbf{h}_{T}^{L}]$ denotes the hidden states of the last layer.
+With $\mathbf{H}^{L}$, a linear projection layer $W$ is introduced to derive the predicted probability distribution $\mathbf{p}_{v}$ over the vocabulary, where $\mathbf{p}_{v}=[\mathbf{p}^{x_{1}}_{v},\mathbf{p}^{x_{2}}_{v},\ldots,\mathbf{p}^{x_{T}}_{v}]$.
+The Transformer model consists of $L$ Transformer layers.
+Each layer is composed of a self-attention sublayer and an FFN sublayer, where FFN is a two-layer fully connected network.
+$\mathbf{\hat{h}}_{1}^{l},\mathbf{\hat{h}}_{2}^{l},\ldots,\mathbf{\hat{h}}_{T}^{l}=SelfAttn(\mathbf{h}_{1}^{l-1},\mathbf{h}_{2}^{l-1},\ldots,\mathbf{h}_{T}^{l-1})$ (4)
+$\mathbf{h}_{1}^{l},\mathbf{h}_{2}^{l},\ldots,\mathbf{h}_{T}^{l}=FFN(\mathbf{\hat{h}}_{1}^{l},\mathbf{\hat{h}}_{2}^{l},\ldots,\mathbf{\hat{h}}_{T}^{l})$ (5)
+$l$ represents the $l$-th transformer layer.
+MoE methods replace the FFN sub-layer with the MoE module.
+A MoE model can replace the FFN with a MoE module for either a single layer, several layers, or even all layers.
+Each MoE module consists of multiple FFNs, whose outputs are mixed with the gating function $P(\cdot)$, represented by the following formula.
+$\mathbf{h}_{t}^{l}=\sum_{i}^{N}P_{i}(\mathbf{\hat{h}}_{t}^{l})\cdot FFN_{i}(\mathbf{\hat{h}}_{t}^{l})$ (6)
+$s.t.|P(\mathbf{\hat{h}}_{t}^{l})|_{0}<<N$ (7)
+Where $N$ represents the number of experts in a single MoE layer.
+The majority of elements in the distribution of a gated $P(\cdot)$ output are zeros, i.e., only a small portion of experts will be activated.
+Therefore, increasing the total number of experts in MoE does not increase computation or inference time.
 3.2 Proposed MaskMoE
-The typical dynamic routing method routes tokens to the top experts based on the scores output by a softmax layer.
-This approach to update expert parameters is sparse, meaning that the parameters of unselected experts are not updated.
-This sparsity may lead to experts underfitting for infrequent tokens.
-Fixed routing ensures that specific experts consistently process the same input tokens, which helps mitigate the problem of insufficient learning for infrequent tokens.
-However, it is less effective in fostering a diversity of representations compared to dynamic routing methods.
-Dynamic routing allows the same token to be encoded by different experts, leading to more varied representations He et al.
-(2023); Yang et al.
-(2019).
-Since frequent tokens can have different meanings depending on the context, fixed routing is less effective than dynamic routing in capturing the nuances of these tokens.
-To balance the density of parameter updates and the diversity of representations in MoE models, we propose MaskMoE to boost token-level learning via routing mask.
-As illustrated in Figure 1, MaskMoE achieves this by introducing a routing masking mechanism.
+Dynamic routing uses a learnable method where tokens are assigned to different experts during training, leading to routing fluctuations.
+On the other hand, fixed routing, like hash layers, assigns each token to a specific expert, which can also limit representation diversity.
+To mitigate the routing fluctuations and the diversity of representations in MoE models, we propose MaskMoE to boost token-level learning via routing mask, as illustrated in Figure 1.
 For each token, a fixed random masking is applied to the experts, allowing them to be routed only to their respective subsets of experts.
 The formula is represented as follows.
-$\mathbf{p}$ $=\text{Softmax}(W_{g}\mathbf{\bar{h}_{t}^{l}}+\mathbf{M})$ (5)
+$\mathbf{p}$ $=\text{Softmax}(W_{g}\mathbf{\hat{h}_{t}^{l}}+\mathbf{m})$ (8)
 Where $\mathbf{m}$ is the masking vector used to control the visibility of token-specific experts.
@@
 For multi-layer MoE models, we reuse the same masking vector $\mathbf{m}$ across different MoE layers.
-Here, $\mathbf{W_{g}}$ is the parameter of the gate.
+Here, $W_{g}$ is the parameter of the gating function.
 The initial energy masking vector for a token $t$ can be represented by the following formula.
@@
 $\mathbf{m}_{j}^{t}$ $=0,\quad\forall j\in\mathbf{C}$
-Formally, we denote the number of visible experts as $V$.
+We denote the number of visible experts as $V$.
+$N$ represents the number of experts in a single MoE layer.
 $\mathcal{U}(\{0:N-1\})$ denotes a uniform distribution over the set of integers from 00 to $N-1$.
+$\text{C}_{i}$ denotes the index of the selected expert.
 It is worth noting that for infrequent tokens, each token has only one visible expert, i.e., $V=1$.
 This promotes more thorough learning by the selected expert for such infrequent tokens.
-For frequent tokens, we also reduce the number of visible experts, but the number of visible experts is greater than 1, i.e., $1<V<=N$.
-For frequent tokens, it is common for $V$ to be less than $N$.
-We believe that an appropriately sized $V$ not only allows frequent words to have higher representation diversity compared to fixed routing but also enables more thorough training for the same token compared to dynamic routing.
+For frequent tokens, the number of visible experts is greater than 1, i.e., $1<V<=N$, and note that it is common for $V$ to be less than $N$.
+We believe that an appropriately sized $V$ not only allows frequent tokens to have higher representation diversity compared to fixed routing but also enables more thorough training for an identical token compared to dynamic routing.
+The complete process of MaskMoE is illustrated in Algorithm 1.
+Algorithm 1 Proposed MaskMoE Algorithm
+0:  Training dataset after tokenization $D$, the vocabulary of the tokenizer $T_{v}$, maximal training step number $S$, the number of visible experts $V$, the number of experts $N$
+1:  Initialize the routing mask matrix, $\mathbf{M}\leftarrow-\infty$
+2:  for each $t$ in $T_{v}$ do
+3:     $\mathbf{C}=\{\text{C}_{i}\}_{i=1}^{V}\sim\mathcal{U}(\{0:N-1\})$
+4:     $\mathbf{m}^{t}=\mathbf{M}(t)$ The corresponding column w.r.t $t$
+5:     $\mathbf{m}_{j}^{t}=0,\quad\forall j\in\mathbf{C}$
+6:  end for
+7:  for $i=1$ to $S$ do
+8:     Load an input sequence $\mathbf{x}$ from $D$
+9:     # Load the routing mask for each token in $\mathbf{x}$
+10:     $\mathbf{m}=\mathbf{M}(\mathbf{x})$
+11:     Train the MoE models with $\mathbf{x}$ as the input, and add $\mathbf{m}$ to the router according to Equation 10
+12:  end for
 3.3 Load Balance Loss
-MoE models typically require distributed training, where different experts are deployed across various nodes.
-To avoid load imbalance, where a minority of expert nodes handle the majority of tokens while the majority of expert nodes spend more time idle, this can impact training efficiency.
+MoE models typically require distributed training, where different experts are deployed across various computation nodes (e.g., GPU server).
+Distributed training can lead to load imbalance, where a minority of expert nodes handle the majority of tokens while others remain idle most of the time, and it can negatively impact training efficiency.
 It is generally desirable for the number of tokens processed by different experts to be roughly equal.
-To achieve load balancing, a load balancing loss is commonly introduced in the training of MoE models.
+To achieve that, a load balancing loss is commonly introduced in the training of MoE models.
 We follow previous works Huang et al.
 (2024); Fedus et al.
-(2022) and adopt a widely used loss.
+(2022) and adopt a widely used loss as follows.
 $\mathcal{L}_{bal}$ $=N*\sum_{i=1}^{N}w_{i}\cdot R_{i}$ (10)
@@
 Where $T$ represents the number of tokens in a mini-batch.
-$\mathbf{p}^{j}$ denotes the probability distribution of the routing output for the j-th token, while $\mathbf{p}_{i}^{j}$ represents the probability value of the i-th expert.
+$\mathbf{p}^{j}$ denotes the probability distribution of the routing output for the $j$-th token, while $\mathbf{p}_{i}^{j}$ represents the specific probability value of the $i$-th expert.
 It is noteworthy that for infrequent tokens, which have only one visible expert, the load balancing loss does not apply to them.
+Since infrequent tokens are assigned to experts through a uniform random hashing, the overall load is supposed to be natively balanced, thus having no need for the balancing loss.
 The balancing loss primarily regulates the routing of frequent tokens.
-Additionally, since infrequent tokens are assigned to experts through random hashing, the overall load is relatively balanced.
-Therefore, our experiments find that MaskMoE, compared to SMoE, does not experience a significant decrease in training efficiency.
 Our final loss is a combination of the language model loss and load-balance loss:
 $\mathcal{L}=\mathcal{L}_{lm}+\mathcal{L}_{bal}$ (13)
-4 Expertiments
+4 Experiments
 Model Configuration Params Activated Params Pile PPL
 Standard Transformer Layers=24, Dense 468M 468M 6.95
+Single-MoE-Layer Setup
 SMoE Layers=24, MoE_Layers = 1 1.3B 468M 6.62
@@
 MaskMoE Layers=24, MoE_Layers = 1 1.3B 468M 6.48
+Multi-MoE-Layer Setup
 SMoE Layers=24, MoE_Layers = 12 10B 468M 6.18
@@
 MaskMoE Layers=24, MoE_Layers = 12 10B 468M 6.11
-Table 1: Perplexity results of language modeling.
+Table 1: Perplexity (PPL) results of language modeling.
 4.1 Pre-training Dataset
 Following previous works, we use the Pile dataset Gao et al.
-(2021) for pre-training data.
-The Pile dataset is a large-scale, publicly available corpus for language model pre-training, containing 22 domains and over 825GB of English text data.
+(2021) as pre-training data.
+The Pile dataset is a large-scale, publicly available corpus, containing 22 domains and over 825GB of English text data.
 For our experiments, we use the well-known LLaMA tokenizer for tokenization, with a vocabulary size of $32k$.
@@
 Then, to identify infrequent tokens and frequent tokens, we calculate the frequency of each token in the training set and sort them in descending order of frequency.
-We categorize the top tokens that cover $P$of the dataset as frequent tokens, and the remaining $(1-P)$as infrequent tokens.
-Where $P$ is a tunable hyperparameter.
-4.2 Experimental Setup
-Following  Touvron et al.
-(2023a, b); Huang et al.
-(2024), we adopt the LLaMA architecture with 24 transformer layers and a hidden layer dimension of 1024.
-In line with LLaMA Touvron et al.
-(2023a), we employ the AdamW  Loshchilov and Hutter (2019) optimizer, and applying a cosine learning rate decay schedule.
-For the MoE model, following  Roller et al.
-(2021), we conduct experiments under both single-layer and multi-layer settings.
-For the single-layer setting, we replace the final FFN layer with the MoE layer.
-For the multi-layer MoE models, we follow Gshard Lepikhin et al.
-(2021), alternating MoE layers every other layer, resulting in a total of $12$ MoE layers in this setting.
-For single-layer MoE models and dense models, we use a learning rate of $43e-$ 4, while for $12$-layer MoE models, we employ a learning rate of $41e-$ 4 to ensure stable convergence of the models.
-If not specified otherwise, tokens that account for the top 40% of frequency are considered frequent tokens, while the remaining are considered infrequent tokens, i.e., $P=0.4$.
-For frequent tokens, the visible expert count $V$ is set to 8, and for infrequent tokens, the visible expert count $V$ is set to 1.
-All of our implementations are based on the DeepSpeed [^1] https://github.com/microsoft/DeepSpeed library Rajbhandari et al.
-(2022); Rasley et al.
-(2020), which offers robust support for MoE-distributed training.
-By default, we enable the random token selection strategy Kim et al.
-(2021) implemented in the library to facilitate faster model convergence and better runtime efficiency.
-Following previous works Roller et al.
-(2021); Fedus et al.
-(2022); Dai et al.
-(2022), unless otherwise specified, our experiments only select the top expert.
+We categorize the top tokens that cover $P\times 100\%$ of the dataset as frequent ones, and the remaining $(1-P)\times 100\%$ as infrequent ones, where $P$ is a tunable hyperparameter, ranging from 0 to 1.
 4.2 Compared Models
-Including MaskMoE, we compare five models for validation experiments. Next, we will briefly introduce five compared models.
-  • Dense represents a standard Transformer language model with 468M parameters.
+We compare the proposed MaskMoE with four models for validation experiments.
+  • Dense represents a standard Transformer language model.
   • SMoE denotes a Switch Transformer Fedus et al. (2022), where the router is a learnable parameter. Unless otherwise specified, each MoE layer has 64 experts.
-  • Hash Layer Roller et al.
-(2021) signifies that through a random hash method, each token is assigned a fixed routing expert prior to training.
-Unless otherwise specified, each MoE layer has 64 experts.
+  • Hash Layer Roller et al. (2021) signifies that through a random hash method, each token is assigned a fixed expert prior to training. Unless otherwise specified, each MoE layer has 64 experts.
   • Share-MoE is a hybrid dense and MoE model created using residual connections Rajbhandari et al.
@@
 This allows for a fair comparison between Share-MoE, SMoE, and Hash Layer.
-  • MaskMoE achieves a balance between representation diversity in MoE layers and thorough parameter training through the routing masking method we proposed.
-Similarly, in this section, MaskMoE also employs the shared expert architecture, which is identical to that of Share-MoE.
+  • MaskMoE aims to achieve a balance between representation diversity in MoE layers and thorough parameter training, via the proposed routing masking method.
+Similarly to Share-MoE, MaskMoE also employs the remarkable shared expert architecture, which has been proven by  Dai et al.
+(2024); Rajbhandari et al.
+(2022) to be beneficial for the training of MoE models.
+4.3 Experimental Setup
+Following  Touvron et al.
+(2023a, b); Huang et al.
+(2024), we adopt the LLaMA architecture with 24 transformer layers and a hidden layer dimension of 1024 and we employ the AdamW  Loshchilov and Hutter (2019) optimizer with a cosine learning rate decay schedule.
+For the MoE model, following  Roller et al.
+(2021), we conduct experiments under both single-layer and multi-layer settings.
+For the single-layer setting, we replace the final FFN sublayer with the MoE sublayer.
+For the multi-layer MoE models, we follow Gshard Lepikhin et al.
+(2021), and the FFN sublayer with MoE sublayer for MoE layers every other layer, resulting in a total of $12$ MoE layers in this setting.
+For single-layer MoE models and dense models, we use a learning rate of $3e^{-4}$, while for $12$-layer MoE models, we employ a learning rate of $1e^{-4}$ to ensure stable convergence of the models.
+If not specified otherwise, tokens that account for the top 40% of frequency are considered frequent tokens, while the remaining are considered infrequent tokens, i.e., $P=0.4$.
+Empirically, for frequent tokens, the visible expert count $V$ is set to 8, and for infrequent tokens, the visible expert count $V$ is set to 1.
+All of our implementations are based on the DeepSpeed [^1] https://github.com/microsoft/DeepSpeed library Rajbhandari et al.
+(2022); Rasley et al.
+(2020), which offers robust support for MoE-distributed training.
+By default, we enable the random token selection method Kim et al.
+(2021) implemented in the library to facilitate faster model convergence and better runtime efficiency.
+Following previous works Roller et al.
+(2021); Fedus et al.
+(2022); Dai et al.
+(2022), unless otherwise specified, our experiments only select the top expert.
 Model BoolQ Hellaswag LAMBADA PIQA SIQA StoryCloze Arc-e TriviaQA WebQs
 Standard Transformer 56.02 40.73 52.55 67.62 40.79 63.55 51.43 7.44 4.97
-MoE_Layers = 1
+Single-MoE-Layer Setup
 SMoE 55.26 43.66 53.46 68.28 41.15 64.14 51.98 10.00 6.99
@@
 MaskMoE 58.38 44.47 55.36 68.99 41.86 65.15 53.14 10.39 7.14
-MoE_Layers = 12
+Multi-MoE-Layer Setup
 SMoE 56.18 46.92 56.37 69.91 41.20 64.99 54.34 13.54 7.33
@@
 MaskMoE 58.32 47.46 57.46 70.62 41.91 65.69 55.35 15.20 6.74
-Table 2: Performances of language models on downstream benchmarks. The best score is marked in bold, and the second best is underlined.
+Table 2: Performances of language models on downstream tasks. The best score is marked in bold, and the second best is underlined.
 4.4 Main Results
-In this section, we first present the model’s perplexity (PPL) on the Pile validation set.
+In this section, we first present the model’s PPL on the Pile validation set.
 Then, following Touvron et al.
@@
 (2024); Dai et al.
-(2024), we conduct tests on a large number of downstream benchmarks, including zero-shot for BoolQ Clark et al.
+(2024), we conduct tests on a large number of downstream benchmarks, including zero-shot tests for BoolQ Clark et al.
 (2019), HellaSwag Zellers et al.
@@
 (2016), and Arc-e Bhakthavatsalam et al.
-(2021), as well as 5-shot for TriviaQA Joshi et al.
+(2021), as well as 5-shot tests for TriviaQA Joshi et al.
 (2017) and WebQs Berant et al.
@@
 Table 1 shows the main results of language modeling on the Pile validation sets.
-With the same parameter activations maintained, MoE models consistently achieve improvements compared to the dense model.
-In comparison to SMoE and Hash Layer, under a single MoE layer setup, MaskMoE has reduced the PPL on the Pile Validation by 0.14 and 0.08 respectively.
-In a 12-MoE layers setup, the PPL reductions are 0.07 and 0.05 respectively.
-Compared to Share-MoE, MaskMoE decreases by 0.24 in single MoE layers and by 0.04 in 12-MoE layers.
-These results demonstrate the rationality of the MaskMoE design.
-It is noteworthy that Share-MoE achieves worse results than SMoE and Hash Layer in the single MoE setup.
+With the same number of activated parameters during inference (i.e., 468M), MoE models consistently achieve improvements over the dense model.
+In comparison to SMoE and Hash Layer, under the Single-MoE-Layer setup, MaskMoE has respectively reduced the PPL on the Pile Validation by 0.14 and 0.08.
+In the Multi-MoE-Layer setup, their corresponding PPL are, respectively, reduced by 0.07 and 0.05.
+Compared to Share-MoE, MaskMoE achieves a PPL decrease of 0.24 in Single-MoE-Layer setup and another PPL decrease of 0.04 in Multi-MoE-Layer setup.
+Such results demonstrate the reasonableness of the MaskMoE design.
+It is noteworthy that Share-MoE achieves worse results than SMoE and Hash Layer in the Single-MoE-Layer setup.
 However, MaskMoE consistently maintains superior performance, whether in single or multi-MoE layers.
-This indicates that MaskMoE exhibits stronger robustness across various settings.
+It indicates that MaskMoE exhibits stronger robustness across various settings.
 4.4.2 Benchmark Results
 As shown in Table 2, we present the performance of the models on downstream tasks.
-Firstly, we observe that MoE models significantly outperform Dense models in downstream tasks.
-Then, we find that MaskMoE substantially outperforms the other MoE baselines, whether in single-layer or 12-layers settings.
-Compared to SMoE, MaskMoE outperforms in all 9 benchmarks with a single MoE layer and in 8 out of 9 benchmarks with 12-MoE layers.
-Against Hash Layer, MaskMoE excels in 7 out of 9 benchmarks with a single MoE layer and in 8 out of 9 with 12-MoE layers.
-Compared to Share-MoE, MaskMoE shows significant improvements in 7 out of 9 benchmarks with a single MoE layer and in 6 out of 9 with 12-MoE layers.
-This indicates that the performance enhancement is achieved by using our proposed router mask to control the visibility of experts for different tokens.
+Firstly, we observe that MoE models consistently and significantly outperform the dense models in downstream tasks.
+And more importantly, we find that MaskMoE substantially outperforms the other MoE baselines, whether in Single-MoE-Layer setup or Multi-MoE-Layer setup.
+Compared to SMoE, MaskMoE consistently outperforms it in all 9 benchmarks with the Single-MoE-Layer setup and 8 out of 9 benchmarks with the Multi-MoE-Layer setup.
+Against Hash Layer, MaskMoE excels in 7 out of 9 benchmarks with the Single-MoE-Layer setup and in 8 out of 9 benchmarks with the Multi-MoE-Layer setup.
+Compared to Share-MoE, MaskMoE demonstrates significant improvements in 7 out of 9 benchmarks with the Single-MoE-Layer setup and in 6 out of 9 benchmarks with the Multi-MoE-Layer setup.
+The result above well demonstrates the performance enhancement achieved by using our proposed router masking method to control the visibility of experts for different tokens.
 We believe the improvements are due to two factors.
-Firstly, compared to SMoE and Share-MoE, infrequent tokens are only routed to a single expert, ensuring more thorough training for these tokens.
-Secondly, in contrast to the Hash Layer, frequent tokens have multiple expert options, which promotes diversity in the representation of frequent tokens.
+Firstly, compared to SMoE and Share-MoE, infrequent tokens are only routed to a single expert, ensuring more thorough training for those tokens.
+Secondly, in contrast to the Hash Layer, frequent tokens have multiple expert options, which promotes diversity in the representation learning of frequent tokens.
 5 Analyses
 We conduct further experiments to provide more insightful analyses on the proposed MaskMoE.
-Considering the constraints of computational resources, following previous works Roller et al.
+Considering the constraints of computational resources and following previous works Roller et al.
 (2021); Xie et al.
@@
 (2022), unless otherwise specified, most of the analytical experiments are conducted under the setting of a single MoE layer.
+5.1 Impact of Shared Expert
+To explore the impact of shared experts, we conduct ablation experiments on them (i.e., removing the shared experts), and the results are shown in Table 3.
+We separately report the average scores [^3] Due to the limited space, here we simply average all metrics on downstream tasks to get an average score for each model.
+of the benchmarks and the PPL of Pile Validation.
+Note that after removing the shared experts, we adopt the same architecture as SMoE and Hash Layer, where the MoE layer consists of 64 fully-sized FFN.
+We find that even without shared experts, MaskMoE still outperforms SMoE, Hash Layer, and Share-MoE in average benchmark scores and PPL.
+Such a result further validates the soundness of our proposed MaskMoE.
+Moreover, incorporating shared expert structures continuously improves MaskMoE’s performance.
+We attribute it to that the shared expert structure promotes greater specialization in token representation DeepSeek-AI et al.
+(2024).
 5.2 Impact of Router Mask
+Model Avg($\uparrow$) PPL($\downarrow$)
+SMoE 43.87 6.62
+Hash Layer 44.56 6.56
+Share-MoE 44.25 6.72
+MaskMoE 44.99 6.48
+w/o Shared Experts 44.85 6.51
+Table 3: Average scores benchmarks and PPL on the Pile validation sets. The best score is marked in bold, and the second best is underlined.
+Considering the proposed MaskMoE without shared experts still outperforms the compared MoE methods, to show the impact of router mask more significantly and minimize the impact of the shared expert, the experiments remove the shared expert.
+Instead, we adopt the same structure as SMoE and Hash Layer, with each MoE layer comprising 64 full-sized FNN modules as experts.
+The performance of MaskMoE is predominantly influenced by two hyperparameters: the boundary threshold for the separation of frequent and infrequent tokens, and the maximum number of visible experts (i.e., $V$) for tokens.
+As shown in Table 4, we conduct an experimental parameter search for both and report their PPL values on the Pile validation sets.
+$V_{a}$ represents the number of experts visible for frequent tokens, and $V_{b}$ represents the number of experts visible for infrequent tokens.
+Notably, when $V_{a}=V_{b}=1$, MaskMoE is equivalent to the Hash Layer, and when $V_{a}=V_{b}=64$, MaskMoE is equivalent to SMoE.
+Then we have the following observations.
+Firstly, we find that MaskMoE outperforms the Hash Layer.
+With $V_{a}=8$ and $V_{b}=1$, MaskMoE reduces ppl by 0.052 on the Pile validation set compared to the Hash Layer, i.e., $V_{a}=V_{b}=1$.
+The result indicates that the multiple-expert design in the proposed MaskMoE for frequent tokens does enhance their representation diversity, which is beneficial for model performance.
+Secondly, compared to SMoE, concentrating tokens into fewer visible experts, such as settings like $V_{a}=V_{b}=8$ or $V_{a}=V_{b}=4$, yields better outcomes.
+It indicates that routing tokens more densely to a reduced number of visible experts can enhance thorough training and improve model performance.
+Moreover, we observe that the configuration $V_{a}=8$ and $V_{b}=1$ outperforms $V_{a}=V_{b}=8$.
+It well verifies that further reducing the number of visible experts is more beneficial for infrequent tokens.
+Finally, it is noteworthy that for MaskMoE, when $V_{b}=1$, regardless of the value of $V_{a}$ (whether it is 4, 8, 16, or 64), the performance of MaskMoE consistently outperforms that of SMoE and Hash Layer.
+That also demonstrates the robustness of MaskMoE.
 $P$ $V_{a}$ $V_{b}$ PPL $P$ $V_{a}$ $V_{b}$ PPL
@@
 / 4 4 6.551 / 8 8 6.566
-Table 3: Impact of Router Mask.
-$\spadesuit$ denotes SMoE, $\clubsuit$ denotes Hash Layer.
-To make the differences in MaskMoE under various configurations more apparent, the values shown in the table are precise to three decimal places.
-To visually demonstrate the impact of the routing masking mechanism, the experiments in this subsection do not employ a shared expert structure.
-Instead, they adopt the same structure as SMoE and Hash Layer, with each MoE layer comprising 64 experts, and the FNN layer is full-sized.
-The performance of MaskMoE is predominantly influenced by two hyperparameters: the boundary threshold $P$ for the division of frequent and infrequent tokens, and the maximum number of visible experts $V$ for tokens.
-As shown in Table 3, we conduct an experimental parameter search for both and reported the Pile Validation PPL.
-$V_{a}$ represents the number of experts visible for frequent tokens, and $V_{b}$ represents the number of experts visible for infrequent tokens.
-Notably, when $V_{a}=V_{b}=1$, MaskMoE is equivalent to the Hash Layer, and when $V_{a}=V_{b}=64$, MaskMoE is equivalent to SMoE.
-We have the following observations.
-Firstly, we find that MaskMoE outperforms the Hash Layer.
-With the $V_{a}=8$ and $V_{b}=1$ settings, MaskMoE reduces perplexity by 0.052 on the Pile validation set compared to the Hash Layer, i.e., $V_{a}=V_{b}=1$.
-This indicates that MaskMoE’s use of multiple expert encodings for frequent tokens enhances representation diversity, which is beneficial for model performance.
-Secondly, compared to SMoE, concentrating tokens into fewer visible experts, such as settings like $V_{a}=V_{b}=8$ or $V_{a}=V_{b}=4$, yields better outcomes.
-This indicates that routing tokens more densely to a reduced number of visible experts enhances thorough training and improves model performance.
-Moreover, we observe that the configuration $V_{a}=8$ and $V_{b}=1$ outperforms $V_{a}=V_{b}=8$.
-This indicates that further reducing the number of visible experts is more beneficial for infrequent tokens.
-Lastly, it is noteworthy that for MaskMoE, when $V_{b}=1$, regardless of the value of $V_{a}$ (whether it is 4, 8, 16, or 64), the performance of MaskMoE consistently outperforms that of SMoE and Hash Layer.
-This demonstrates the robustness of MaskMoE.
-5.2 Impact of Shared Expert
-To explore the impact of shared experts, in this subsection, we conduct ablation experiments on shared experts, and the results are shown in Table 4.
-We separately report the average scores of the benchmarks and the PPL of Pile Validation.
-It is noteworthy that after removing the shared experts, we adopt the same architecture as SMoE and Hash Layer, where the MoE layer consists of 64 fully-sized FFN.
-We find that even without shared experts, MaskMoE outperforms SMoE, Hash Layer, and Share-MoE in average benchmark scores and Pile Validation PPL.
-This further validates the soundness of MaskMoE’s design.
-Additionally, we observe that incorporating shared expert structures continuously improves MaskMoE’s performance.
-We believe this is because the shared expert structure promotes greater specialization in token representation DeepSeek-AI et al.
-(2024).
-Model Avg($\uparrow$) PPL($\downarrow$)
-SMoE 43.87 6.62
-Hash Layer 44.56 6.56
-Share-MoE 44.25 6.72
-MaskMoE 44.99 6.48
-w/o Shared Experts 44.85 6.51
-Table 4: Average scores of benchmarks and perplexity on the Pile validation. The best score is marked in bold, and the second best is underlined.
+Table 4: Impact of Router Mask. $\spadesuit$ denotes SMoE, $\clubsuit$ denotes Hash Layer.
 5.3 Number of Experts
-(a) Non-Shared Expert Structures
-(b) Shared Expert Structures
+To investigate the variation in model performance under different configurations of expert numbers, we conducted experiments with 32, 64, and 128 experts.
+The experimental results, as shown in Figure 2, reveal the following observations.
+First of all, we observe that for SMoE and Share-MoE, increasing the number of experts does not necessarily lead to a lower PPL.
+In fact, when the number of experts reaches 128, both models even exhibit higher PPLs compared to having only 64 experts.
+We attribute it to that having more experts leads to the occurrences of an identical token being spread among more experts, resulting in each expert not learning enough about this token, which leads to model underfitting.
+In contrast, the performance of Hash Layer and MaskMoE continue to improve with the increase in the number of experts.
+We attribute it to that, despite the growing number of experts, the visible experts for an identical token in Hash Layer and MaskMoE remain a fixed, small number, allowing the experts to continue learning adequately.
+It indicates that they are more favorable for scaling the number of experts.
+Secondly, we observe that compared to the Hash Layer, MaskMoE consistently performs better across all configurations.
+Although the Hash Layer also enables experts to be adequately trained with the tokens assigned to them, MaskMoE excels in the diversity of representations learning for frequent tokens.
+such a superiority in handling token diversity contributes to MaskMoE’s overall advantage over Hash Layer.
+Finally, we find that regardless of the number of experts, MaskMoE consistently maintains the best performance.
+That further verifies the superiority of the proposed MaskMoE and the reasonableness of its design.
+(a) None-Shared Expert Structures
+(b) Shared-Expert Structures
 Figure 2: Comparison of MoE-based Transformers with different numbers of experts.
-To investigate the variation in model performance under different configurations of expert numbers, we conducted experiments with 16, 64, and 128 experts.
-The experimental results, as shown in Figure 2, reveal the following observations.
-Firstly, we observe that for SMoE and Share-MoE, increasing the number of experts does not necessarily lead to a lower perplexity.
-In fact, when the number of experts reaches 128, the model exhibits a higher PPL compared to having only 64 experts.
-We believe this is because with more experts, the average number of tokens assigned to each expert decreases, leading to further underfitting of the model.
-In contrast, the performance of the Hash Layer and MaskMoE consistently improves with the increase in the number of experts, indicating that they are more favorable for scaling the number of experts.
-Secondly, we observe that compared to the Hash Layer, MaskMoE consistently performs better across all configurations.
-Although the Hash Layer also enables experts to be adequately trained with the tokens assigned to them, MaskMoE excels in the diversity of representations for frequent tokens.
-This superiority in handling token diversity contributes to MaskMoE’s overall advantage over Hash Layer.
-Lastly, we find that regardless of the number of experts, MaskMoE consistently maintains the best performance. This indicates that MaskMoE has strong parameter robustness.
 6 Conclusions
-In this paper, we analyze dynamic routing models such as SMoE, which routes the same token to multiple experts, potentially causing the model to underfit, especially for infrequent tokens.
-While fixed routing can mitigate this issue, it struggles with capturing rich features for frequent tokens.
-To address these issues, we propose a routing masking method aimed at ensuring adequate learning of infrequent tokens while maintaining rich representations for frequent tokens.
-Extensive experiments demonstrate that the proposed MaskMoE outperforms various downstream tasks and significantly improves perplexity (PPL) on the Pile dataset compared to SMoE, Hash Layer, and Share-MoE.
+In existing MoE language models, the widely-used dynamic routing models such as SMoE, which routes the same token to multiple experts, potentially cause the model to encounter underfitting during training, especially for infrequent tokens.
+Although fixed routing like Hash Layer can mitigate the routing fluctuation issue, it struggles with capturing rich features for frequent tokens.
+To tackle the routine fluctuation issue and the representation diversity learning issue at the meantime, we propose a routing masking method termed MaskMoE, aiming at ensuring adequate learning of infrequent tokens while maintaining rich representations learning for frequent tokens.
+Extensive experiments demonstrate that the proposed MaskMoE outperforms existing MoE methods like SMoE, Hash Layer, and Share-MoE on various downstream tasks and significantly reduces PPL on the Pile validation sets.
+That will demonstrates the effectiveness of MaskMoE.
 7 Limitiations
 In this paper, we categorize tokens simply into frequent and infrequent tokens by their frequency.
-This classification seems somewhat rigid.
-Considering a smoother classification method could potentially yield better results.
-Specifically, when tokens are arranged in ascending order, their frequencies demonstrate a gradual decline, and the corresponding number of visible experts also decreases progressively.
+Such a classification method seems somewhat rigid.
+A smoother classification method could potentially yield better results for our proposed MaskMoE.
 Due to computational constraints, we do not conduct further experiments.
-We leave the exploration of smoother partitioning methods for future work.
+We leave the exploration of smoother partitioning methods for our future work.

### v4

@@
 Mixture-of-Experts models (MoE) address the issue by allowing model size to scale up without substantially increasing training or inference costs.
+In MoE, there is an important module called the router, which is used to distribute each token to the experts.
+Currently, the mainstream routing methods include dynamic routing and fixed routing.
 Despite their promising results, MoE models encounter several challenges.
 Primarily, for dynamic routing methods, the dispersion of training tokens across multiple experts can lead to underfitting, particularly for infrequent tokens.
-Additionally, while fixed routing methods can mitigate that issue, they compromise on the diversity of representations.
+Additionally, though fixed routing methods can mitigate that issue, they compromise on the diversity of representations.
 In this paper, we propose MaskMoE, a method designed to enhance token-level learning by employing a routing masking technique within the Mixture-of-Experts model.
@@
 (2020); Anil et al.
+(2023); Chowdhery et al.
+(2022); Radford et al.
+(2019); Rae et al.
+(2021); Biderman et al.
 (2023).
@@
 (2020); Hoffmann et al.
-(2022), as the model size increases, the model’s size will continue to grow.
-However, for dense language models, the computational costs of continuing to scale up are substantial.
+(2022), as the model size increases, the model’s capabilities will continue to grow.
+However, for dense language models, the computational costs of continuing to scale up are excessive.
 In order to further scale up models within computational budgets, sparse activation networks Child et al.
@@
 (2022); Roller et al.
-(2021), which trains multiple expert networks but selects only a subset to process specific inputs.
-Compared to dense networks of the same model size, MoE effectively reduces computational costs and achieves better results in both PPL and downstream tasks Lepikhin et al.
+(2021), which trains multiple expert layers but selects only a subset to process specific inputs.
+Compared to dense networks of the same model size, MoE effectively reduces computational costs and achieves comparable results in both PPL and downstream Lepikhin et al.
 (2021); Du et al.
@@
 (2024).
-In MoE models, for each input (e.g., a token in the language models), the router in front of the experts needs to decide which tokens to distribute to which experts.
-The current commonly used dynamic routing method in MoE is to select the expert with the top $k$ highest confidence based on a probability distribution output by an intermediate layer with learnable parameters.
+In MoE models, for each input (e.g., a token in the language models), the router in front of the experts needs to decide which experts to feed it in.
+The commonly used dynamic routing method in MoE is to select the expert with the top $k$ highest confidence based on a probability distribution output by an intermediate layer with learnable parameters that act as the router.
 Previous works Lepikhin et al.
@@
 (2022) show that MoE models trained based on dynamic routing achieve better performance than dense models with the same amount of training computation.
-However, this dynamic routing method introduces a new challenge: the routing fluctuation problem Dai et al.
+Figure 1: Illustration of random routing masking method. For each token, some experts will be randomly masked, with only a subset of experts being visible.
+However, the dynamic routing method introduces a new challenge: the routing fluctuation problem Dai et al.
 (2022).
-It implies that as training moves on, the same token is assigned to different experts in different iterations.
-This has two negative impacts.
-Firstly, each expert is trained on only a subset of the occurrences of an identical token, resulting in each expert not learning enough about this token.
-As previous work Shazeer et al.
+It implies that as training moves on, the same token is assigned to different experts in different iterations due to the variation of learnable parameters in the router.
+That has two negative impacts.
+a) Each expert is trained on only a subset of the occurrences of an identical token, resulting in each expert possibly not learning enough about the token.
+As  Shazeer et al.
 (2017) has indicated, when there are too many experts, the MoE model exhibits underfitting, and even shows higher PPL than its counterparts with fewer experts.
 The routing fluctuation problem has relatively less impact on the learning of frequent tokens, as those tokens have sufficient occurrences to ensure that each expert receives adequate training.
-On the contrary, for infrequent tokens, the routing fluctuation problem can disperse each of them across various experts, which can lead to underfitting for them.
-Secondly, the knowledge each expert learns from the corresponding occurrence subset w.r.t an identical token is hard to share among experts.
-Such lack of knowledge sharing and interaction among experts may negatively impact the model’s performance Zhao et al.
+On the contrary, for infrequent tokens, the routing fluctuation problem can disperse each of them across various experts, which probably leads to underfitting for them.
+b) The knowledge each expert learns from the corresponding occurrence subset w.r.t an identical token is hard to share among experts.
+Then the underfitting of experts for some tokens, together with the lack of knowledge sharing among experts, can negatively impact the model’s performance Zhao et al.
 (2024).
-The above two issues hinder scaling the number of experts for an MoE model.
-Because, as the number of experts increases while keeping the training data constant, the average number of tokens assigned to each expert further decreases, exacerbating the above two issues.
-Figure 1: Illustration of random routing masking method.
-For each token, some experts will be randomly masked, with only a subset of experts being visible, e.g., for the token “in” only expert FFN #1, #2, and #3 are visible.
-To alleviate the issue of routing fluctuations, the previously proposed fixed routing method Roller et al.
+Both issues above hinder scaling the number of experts for an MoE model.
+Because, as the number of experts increases while keeping the training data unchanged, the average number of token occurrences assigned to each expert further decreases, exacerbating the issues.
+To alleviate the issue of routing fluctuations, fixed routing method Roller et al.
 (2021); Dai et al.
-(2022) is a solution.
+(2022) is proposed, e.g., Hash Layer.
 Specifically, Hash Layer Roller et al.
-(2021) pre-assigns the same token to a fixed routing expert.
-For instance, as shown in Figure 1, the token “deliberate” can only be routed to the expert FFN#2.
+(2021) pre-assigns each token to a fixed expert.
+For instance, as shown in Figure 1, all occurrences of the token “deliberate” can only be routed to the expert FFN#2.
 The fixed routing method to send all the occurrences of a token to the same expert is advantageous for the token’s thorough learning, especially for infrequent tokens.
-However, fixed routing based on the Hash Layer represents a deficiency in diversity, especially for those frequent tokens.
+However, fixed routing based on the Hash Layer represents a deficiency in learning representation diversity, especially for those frequent tokens.
 Previous works He et al.
 (2023); Yang et al.
-(2019) show that increasing the number of optional experts for all occurrences of an identical can enhance the diversity of representations.
-Frequent tokens probably require to be encoded by more experts due to their broader range of usage Koranda et al.
+(2019) show that increasing the number of optional experts for all occurrences of an identical token can help the diversity of learned representations.
+And frequent tokens probably require to be encoded by more experts due to their broader range of usage Koranda et al.
 (2018).
-For instance, the token “the” appears in various contexts and owns different semantic meanings, and having a more diverse representation helps to distinguish it.
-Therefore, based on the observations above, we need a fixed routing strategy for infrequent tokens, while also needing more experts for frequent ones.
-To meet the requirements above.
-In this paper, we propose a routing method called MaskMoE, which includes a routing masking method that adjusts the number of visible experts for tokens of different frequencies, by generating a masking vector for each token in the vocabulary before training.
+For instance, the token “play” appears in various contexts and owns different semantic meanings, and thus having a more diverse representation helps to distinguish them.
+Therefore, based on the observations above, we need a fixed routing strategy to alleviate the underfitting problem for infrequent tokens, while also requiring more experts for frequent ones to maintain the representational diversity.
+To meet the requirements above, in this paper, we propose a routing method called MaskMoE, which includes a routing masking method that adjusts the number of visible experts for tokens of different frequencies, by generating a masking vector for each token in the vocabulary before training.
 Specifically, for each infrequent token, the proposed MaskMoE employs routing masking to retain only one visible expert to which the token can be routed, in each MoE layer.
-For each frequent token, the proposed MaskMoE has a lower routing masking rate to allow more visible experts for it in each MoE layer.
+For each frequent token, the proposed MaskMoE allow more visible experts for it in each MoE layer.
 MaskMoE can enhance token-level learning through such a routing mask.
-Specifically, MaskMoE makes the model learn more intensively about infrequent tokens by routing each to an identical expert in each MoE layer, enabling the expert to be fully trained w.r.t that token.
+Specifically, MaskMoE makes the model learn more intensively about infrequent tokens by routing each to an identical expert in every MoE layer, enabling the expert to be adequately trained w.r.t that token.
 Meanwhile, MaskMoE maintains the diversity of representations for frequent tokens by still having multiple experts available for routing, which benefits their representation diversity He et al.
@@
   • We highlight that dynamic routing, which disperses the occurrences of a token into different experts during training, can lead to the underfitting of experts on infrequent tokens, while fixed routing lacks representation diversity for frequent tokens.
-  • We validate the effectiveness of the proposed MaskMoE with extensive experiments. Experimental results show that it consistently outperforms previous Mixture-of-Experts models.
+  • We validate the effectiveness of the proposed MaskMoE with extensive experiments.
+Experimental results show that it consistently outperforms previous dynamic routing and fixed routing methods for MoE models.
 2 Related Work
@@
 (2023a, b).
-Currently, GPT4 OpenAI (2023), trained with Reinforcement Learning from Human Feedback (RLHF), achieves truly remarkable results.
+Currently, GPT4 OpenAI (2023) achieves truly remarkable results.
 However, as the size of the model grows, the computational demands for both training and inference also increase.
@@
 3 Method
-In this section, we first review the training of the MoE models. Then, we elaborate on the proposed MaskMoE. Finally, we introduce the load balance loss.
+3.1 Reviewing the Language Models
+Given a tokenized input sequence $\mathbf{x}=({x_{1},x_{2},...,x_{T}})$ consisting of $T$ tokens, a language model generates a probability distribution $\mathbf{p}$ over the vocabulary as output for each token.
+In nearly all implementations, the Cross-Entropy loss is used as the loss function to maximize the predicted probability $\mathbf{P}_{\cdot,i}^{x_{i}}$ w.r.t the ground-truth token $x_{i}$.
+The training loss $\mathcal{L}_{lm}$ of the generative language model can be formulated as follows:
+$\mathcal{L}_{lm}$ $=-\sum_{i=2}^{T}\log(\mathbf{P}_{\cdot,i}^{x_{i}})$ (1)
+$s.t.,\quad\mathbf{P}_{\cdot,i}$ $=\text{softmax}(W\mathbf{H}^{L}_{\cdot,i})$
+$\mathbf{H}^{L}$ $=\text{Transformer}(x_{1},x_{2},\ldots,x_{T})$
+where $\mathbf{P}_{\cdot,i}$ and $\mathbf{H}_{\cdot,i}^{L}$ denote the $i$-th column of the matrix $\mathbf{P}$ and $\mathbf{H}^{L}$.
+Here, $\mathbf{H}^{L}=[\mathbf{h}_{1}^{L},\mathbf{h}_{2}^{L},\ldots,\mathbf{h}_{T}^{L}]$ denotes the hidden states of the last layer.
+With $\mathbf{H}^{L}_{\cdot,i}$, a linear projection layer $W$ is introduced to derive the predicted probability distribution $\mathbf{P}_{\cdot,i}$ over the vocabulary.
+The Transformer model consists of $L$ Transformer blocks.
+Each block is composed of a multi-head self-attention (MHA) module and an FFN module, where the FFN module is generally a two-layer fully connected network.
+Formally,
+$\mathbf{\hat{h}}_{1}^{l},\mathbf{\hat{h}}_{2}^{l},\ldots,\mathbf{\hat{h}}_{T}^{l}=\text{MHA}(\mathbf{h}_{1}^{l-1},\mathbf{h}_{2}^{l-1},\ldots,\mathbf{h}_{T}^{l-1})$ (2)
+$\mathbf{h}_{1}^{l},\mathbf{h}_{2}^{l},\ldots,\mathbf{h}_{T}^{l}=\text{FFN}(\mathbf{\hat{h}}_{1}^{l},\mathbf{\hat{h}}_{2}^{l},\ldots,\mathbf{\hat{h}}_{T}^{l})$ (3)
+where $l$ represents the $l$-th Transformer block.
 3.2 Reviewing the Mixture-of-Experts
-The routing of MoE can be divided into fixed routing Roller et al.
+MoE methods generally replace a single FFN module or multiple/all FFN modules as MoE modules.
+Each MoE module consists of multiple FFNs, whose outputs are mixed with the routine function $\mathbf{r}(\cdot)$, i.e., the router, as shown in the following formula:
+$\mathbf{h}_{t}^{l}=\sum_{i}^{N}\mathbf{r}_{i}(\mathbf{\hat{h}}_{t}^{l})\cdot\text{FFN}_{i}(\mathbf{\hat{h}}_{t}^{l})\qquad s.t.,|\mathbf{r}(\mathbf{\hat{h}}_{t}^{l})|_{0}\ll N$ (4)
+where $N$ represents the number of experts in a single MoE module, $\mathbf{r}_{i}$ denotes the routine result w.r.t the $i$-th expert, and $|\cdot|_{0}$ denotes the $L_{0}$-norm.
+The majority of elements $\mathbf{r}(\cdot)$ are zeros, and thus only a small portion of experts will be activated.
+Therefore, increasing the total number of experts in MoE does not increase computation or inference time substantially.
+Regarding the router $\mathbf{r}$ of MoE, it can be divided into fixed routing  Roller et al.
 (2021); Dai et al.
-(2022) and dynamic routing with learnable parameters Lepikhin et al.
-(2021); Du et al.
-(2022a); Dai et al.
-(2024); Fedus et al.
+(2022) and dynamic routing with learnable parameters  Fedus et al.
+(2022); Lepikhin et al.
+(2021); DeepSeek-AI et al.
+(2024); Dai et al.
 (2024).
-The common method for fixed routing is to use random hashing to determine Roller et al.
-(2021), before training, which experts each token will be sent to.
-In contrast, dynamic routing employs a routing layer with learnable parameters to decide which expert should process an input token.
-In terms of structure, the FFN layer of a dense language model is generally replaced with sparse expert modules to build an MoE model.
-Given a tokenized input sequence $\mathbf{x}={x_{1},x_{2},...,x_{T}}$ consisting of $T$ tokens , and then generates a probability distribution $\mathbf{p}_{v}$ over the vocabulary as output.
-In nearly all implementations, the Cross-Entropy loss is used as the loss function to maximize the predicted probability $\mathbf{p}_{v}^{x_{i}}$ w.r.t the ground-truth token $x_{i}$.
-The training loss $\mathcal{L}_{lm}$ of the generative language model can be formulated as follows.
-$\mathcal{L}_{lm}$ $=-\sum_{i=2}^{T}\log(\mathbf{p}_{v}^{x_{i}})$ (1)
-$s.t.,\mathbf{p}_{v}$ $=\text{softmax}(W\mathbf{H}^{L})$ (2)
-$\mathbf{H^{L}}$ $=\text{Transformer}(x_{1}^{0},x_{2}^{0},\ldots,x_{T}^{0})$ (3)
-Here, $\mathbf{H^{L}}=[\mathbf{h}_{1}^{L},\mathbf{h}_{2}^{L},\ldots,\mathbf{h}_{T}^{L}]$ denotes the hidden states of the last layer.
-With $\mathbf{H}^{L}$, a linear projection layer $W$ is introduced to derive the predicted probability distribution $\mathbf{p}_{v}$ over the vocabulary, where $\mathbf{p}_{v}=[\mathbf{p}^{x_{1}}_{v},\mathbf{p}^{x_{2}}_{v},\ldots,\mathbf{p}^{x_{T}}_{v}]$.
-The Transformer model consists of $L$ Transformer layers.
-Each layer is composed of a self-attention sublayer and an FFN sublayer, where FFN is a two-layer fully connected network.
-$\mathbf{\hat{h}}_{1}^{l},\mathbf{\hat{h}}_{2}^{l},\ldots,\mathbf{\hat{h}}_{T}^{l}=SelfAttn(\mathbf{h}_{1}^{l-1},\mathbf{h}_{2}^{l-1},\ldots,\mathbf{h}_{T}^{l-1})$ (4)
-$\mathbf{h}_{1}^{l},\mathbf{h}_{2}^{l},\ldots,\mathbf{h}_{T}^{l}=FFN(\mathbf{\hat{h}}_{1}^{l},\mathbf{\hat{h}}_{2}^{l},\ldots,\mathbf{\hat{h}}_{T}^{l})$ (5)
-$l$ represents the $l$-th transformer layer.
-MoE methods replace the FFN sub-layer with the MoE module.
-A MoE model can replace the FFN with a MoE module for either a single layer, several layers, or even all layers.
-Each MoE module consists of multiple FFNs, whose outputs are mixed with the gating function $P(\cdot)$, represented by the following formula.
-$\mathbf{h}_{t}^{l}=\sum_{i}^{N}P_{i}(\mathbf{\hat{h}}_{t}^{l})\cdot FFN_{i}(\mathbf{\hat{h}}_{t}^{l})$ (6)
-$s.t.|P(\mathbf{\hat{h}}_{t}^{l})|_{0}<<N$ (7)
-Where $N$ represents the number of experts in a single MoE layer.
-The majority of elements in the distribution of a gated $P(\cdot)$ output are zeros, i.e., only a small portion of experts will be activated.
-Therefore, increasing the total number of experts in MoE does not increase computation or inference time.
+The former commonly use random hashing to determine  Roller et al.
+(2021), before training, which experts each token will be sent to, i.e., $\mathbf{r}$ is unlearnable and preset.
+In contrast, the latter employs a routing layer with learnable parameters to decide which expert should process an input token, i.e., $\mathbf{r}$ is to be learned during model training.
 3.3 Proposed MaskMoE
-Dynamic routing uses a learnable method where tokens are assigned to different experts during training, leading to routing fluctuations.
-On the other hand, fixed routing, like hash layers, assigns each token to a specific expert, which can also limit representation diversity.
-To mitigate the routing fluctuations and the diversity of representations in MoE models, we propose MaskMoE to boost token-level learning via routing mask, as illustrated in Figure 1.
-For each token, a fixed random masking is applied to the experts, allowing them to be routed only to their respective subsets of experts.
-The formula is represented as follows.
-$\mathbf{p}$ $=\text{Softmax}(W_{g}\mathbf{\hat{h}_{t}^{l}}+\mathbf{m})$ (8)
-Where $\mathbf{m}$ is the masking vector used to control the visibility of token-specific experts.
-For visible experts, the corresponding elements in $\mathbf{m}$ are 0, while for invisible experts, the corresponding elements in $\mathbf{m}$ are $\infty$- ∞.
+As mentioned before, dynamic routing can cause routing fluctuations, while fixed routing, like hash layers, can limit representation diversity.
+To mitigate the above two issues in MoE models, we propose MaskMoE to boost token-level learning via routing mask, as illustrated in Figure 1.
+Specifically, for each token, a fixed random masking is applied to the experts, allowing them to be routed only to their corresponding subsets of experts.
+The formula is represented as follows:
+$\mathbf{r}$ $=\text{softmax}(W_{r}\mathbf{\hat{h}_{t}}^{l}+\mathbf{m}^{t})$ (5)
+where $\mathbf{m}^{t}$ is the masking vector used to control the visibility of token-specific experts for the token $t$.
+For visible experts, the corresponding elements in $\mathbf{m}^{t}$ are 0, while for invisible experts, the corresponding elements in $\mathbf{m}^{t}$ are $\infty$- ∞.
 The masking vector is determined before training and does not change during the training process.
-For multi-layer MoE models, we reuse the same masking vector $\mathbf{m}$ across different MoE layers.
-Here, $W_{g}$ is the parameter of the gating function.
-The initial energy masking vector for a token $t$ can be represented by the following formula.
+For multi-layer MoE models, we reuse the same masking vector $\mathbf{m}^{t}$ across different MoE layers.
+Here, $W_{r}$ is the learnable parameter of the routing function.
+It can be seen that we use the preset $\mathbf{m}^{t}$ to control the visibility of experts like the fixed routing method and meanwhile reserve the learnable parameter $W_{r}$ as the dynamic routing method, and thus we combine the strengths of both.
+The initial masking vector for a token $t$ can be represented by the following formula:
 $\mathbf{m}^{t}$ $:=-\infty\cdot\mathbf{1}_{N}$ (6)
-$\mathbf{C}$ =$=\{\text{C}_{i}\}_{i=1}^{V}\sim\mathcal{U}(\{0:N-1\})$
+$\mathbf{C}$ =$=\{\text{C}_{i}\}_{i=1}^{V}\sim\mathcal{U}(\{1:N\})$
 $\mathbf{m}_{j}^{t}$ $=0,\quad\forall j\in\mathbf{C}$
-We denote the number of visible experts as $V$.
-$N$ represents the number of experts in a single MoE layer.
-$\mathcal{U}(\{0:N-1\})$ denotes a uniform distribution over the set of integers from 00 to $N-1$.
-$\text{C}_{i}$ denotes the index of the selected expert.
-It is worth noting that for infrequent tokens, each token has only one visible expert, i.e., $V=1$.
-This promotes more thorough learning by the selected expert for such infrequent tokens.
-For frequent tokens, the number of visible experts is greater than 1, i.e., $1<V<=N$, and note that it is common for $V$ to be less than $N$.
+where the number of visible experts is $V$.
+$\mathcal{U}(\{1:N\})$ denotes a uniform distribution over the set of integers from $1$ to $N$, and $\text{C}_{i}$ denotes the index of the selected expert for the $i$-th random sampling.
+It is worth noting that for infrequent tokens, each token usually has only one visible expert, i.e., $V=1$.
+That promotes more thorough learning by the selected expert for such infrequent tokens.
+For frequent tokens, the number of visible experts is generally greater than 1, i.e., $1<V<=N$, and it is common for $V$ to be less than $N$.
 We believe that an appropriately sized $V$ not only allows frequent tokens to have higher representation diversity compared to fixed routing but also enables more thorough training for an identical token compared to dynamic routing.
-The complete process of MaskMoE is illustrated in Algorithm 1.
-Algorithm 1 Proposed MaskMoE Algorithm
-0:  Training dataset after tokenization $D$, the vocabulary of the tokenizer $T_{v}$, maximal training step number $S$, the number of visible experts $V$, the number of experts $N$
-1:  Initialize the routing mask matrix, $\mathbf{M}\leftarrow-\infty$
-2:  for each $t$ in $T_{v}$ do
-3:     $\mathbf{C}=\{\text{C}_{i}\}_{i=1}^{V}\sim\mathcal{U}(\{0:N-1\})$
-4:     $\mathbf{m}^{t}=\mathbf{M}(t)$ The corresponding column w.r.t $t$
-5:     $\mathbf{m}_{j}^{t}=0,\quad\forall j\in\mathbf{C}$
-6:  end for
-7:  for $i=1$ to $S$ do
-8:     Load an input sequence $\mathbf{x}$ from $D$
-9:     # Load the routing mask for each token in $\mathbf{x}$
-10:     $\mathbf{m}=\mathbf{M}(\mathbf{x})$
-11:     Train the MoE models with $\mathbf{x}$ as the input, and add $\mathbf{m}$ to the router according to Equation 10
-12:  end for
-3.3 Load Balance Loss
-MoE models typically require distributed training, where different experts are deployed across various computation nodes (e.g., GPU server).
-Distributed training can lead to load imbalance, where a minority of expert nodes handle the majority of tokens while others remain idle most of the time, and it can negatively impact training efficiency.
-It is generally desirable for the number of tokens processed by different experts to be roughly equal.
-To achieve that, a load balancing loss is commonly introduced in the training of MoE models.
-We follow previous works Huang et al.
-(2024); Fedus et al.
-(2022) and adopt a widely used loss as follows.
-$\mathcal{L}_{bal}$ $=N*\sum_{i=1}^{N}w_{i}\cdot R_{i}$ (10)
-[…]
-$R_{i}$ $=\frac{1}{T}\sum_{j=1}^{T}\mathbf{p}_{i}^{j}$ (12)
-Where $T$ represents the number of tokens in a mini-batch.
-$\mathbf{p}^{j}$ denotes the probability distribution of the routing output for the $j$-th token, while $\mathbf{p}_{i}^{j}$ represents the specific probability value of the $i$-th expert.
-It is noteworthy that for infrequent tokens, which have only one visible expert, the load balancing loss does not apply to them.
-Since infrequent tokens are assigned to experts through a uniform random hashing, the overall load is supposed to be natively balanced, thus having no need for the balancing loss.
-The balancing loss primarily regulates the routing of frequent tokens.
-Our final loss is a combination of the language model loss and load-balance loss:
-$\mathcal{L}=\mathcal{L}_{lm}+\mathcal{L}_{bal}$ (13)
-4 Experiments
-Model Configuration Params Activated Params Pile PPL
+Model Configuration Params Activated Params Pile PPL ($\downarrow$)
 Standard Transformer Layers=24, Dense 468M 468M 6.95
@@
 Table 1: Perplexity (PPL) results of language modeling.
+3.4 Load Balance Loss
+LLMs are generally trained in a distributed manner.
+Yet distributed training can lead to load imbalance of MoE models Lepikhin et al.
+(2021); Fedus et al.
+(2022); Dai et al.
+(2022), where a minority of expert handle the majority of tokens while the others remain idle most of the time.
+That can negatively impact training efficiency.
+It is generally desirable for the number of tokens processed by different experts to be roughly equal.
+To achieve that, a load balancing loss is commonly introduced in the training of MoE models.
+We follow previous works Huang et al.
+(2024); Fedus et al.
+(2022) and adopt a widely used loss as follows:
+$\mathcal{L}_{bal}$ $=N\cdot\sum_{i=1}^{N}w_{i}R_{i}$ (7)
+[…]
+$R_{i}$ $=\frac{1}{B}\sum_{j=1}^{B}\mathbf{r}_{i}^{j}$
+where $B$ represents the number of tokens in a mini-batch.
+$\mathbf{r}^{j}$ denotes the probability distribution of the routing output for the $j$-th token, derived via Eq.
+7, and $\mathbf{r}_{i}^{j}$ represents the specific probability value w.r.t the $i$-th expert.
+It is noteworthy that for infrequent tokens, which have only one visible expert, the load balancing loss does not apply to them.
+Since infrequent tokens are assigned to experts through uniform random hashing, the overall load is supposed to be natively balanced, thus having no need for the balancing loss.
+The balancing loss primarily regulates the routing w.r.t frequent tokens.
+Our final loss is a combination of the language model loss and the load-balance loss:
+$\mathcal{L}=\mathcal{L}_{lm}+\mathcal{L}_{bal}$ (8)
+4 Experiments
 4.1 Pre-training Dataset
@@
 (2024) to calculate the sampling rate for each domain based on the number of tokens after tokenization.
-Due to the computational budget, we follow the pretraining settings of  Xie et al.
+Due to the limited computational budget, and following the pretraining settings of  Xie et al.
 (2023a); Su et al.
@@
 Then, to identify infrequent tokens and frequent tokens, we calculate the frequency of each token in the training set and sort them in descending order of frequency.
-We categorize the top tokens that cover $P\times 100\%$ of the dataset as frequent ones, and the remaining $(1-P)\times 100\%$ as infrequent ones, where $P$ is a tunable hyperparameter, ranging from 0 to 1.
+We categorize the top tokens that cover $P\times 100\%$ of the dataset as frequent ones, and the remaining as infrequent ones, where $P$ is a tunable hyperparameter, ranging from 0 to 1.
 4.2 Compared Models
-We compare the proposed MaskMoE with four models for validation experiments.
+We compare the proposed MaskMoE with four remarkable baselines for validation experiments.
+Following previous works Roller et al.
+(2021); Fedus et al.
+(2022); Dai et al.
+(2022), unless otherwise specified, our experiments select the top-$1$ expert for each MoE layer.
   • Dense represents a standard Transformer language model.
-  • SMoE denotes a Switch Transformer Fedus et al. (2022), where the router is a learnable parameter. Unless otherwise specified, each MoE layer has 64 experts.
-  • Hash Layer Roller et al. (2021) signifies that through a random hash method, each token is assigned a fixed expert prior to training. Unless otherwise specified, each MoE layer has 64 experts.
+  • SMoE denotes a Switch Transformer Fedus et al. (2022), where the router is a learnable layer. Unless otherwise specified, each MoE layer has 64 experts.
+  • Hash Layer Roller et al. (2021) signifies that through a random hash method, each token is assigned to a fixed expert before training. Unless otherwise specified, each MoE layer has 64 experts.
   • Share-MoE is a hybrid dense and MoE model created using residual connections Rajbhandari et al.
@@
 (2022).
-Unless otherwise specified, Share-MoE has 1 shared expert and 128 routed experts, where each expert is 0.5 times the size of a standard FFN.
-During both training and inference, in addition to the activation of shared experts, the top expert is also selected from a pool of 128 experts.
+Unless otherwise specified, Share-MoE has 1 shared expert and 128 routed experts, where each expert is 0.5$\times$ of the size of a standard FFN.
+During both training and inference, in addition to the activation of shared experts, the top-$1$ expert is also selected from the pool of 128 experts.
 In such a setup, Share-MoE has the same number of floating-point operations (FLOPs) and activation parameters as SMoE and Hash Layer.
-This allows for a fair comparison between Share-MoE, SMoE, and Hash Layer.
-  • MaskMoE aims to achieve a balance between representation diversity in MoE layers and thorough parameter training, via the proposed routing masking method.
+That allows for a fair comparison between Share-MoE, SMoE, and Hash Layer.
+  • MaskMoE aims to achieve a balance between representation diversity and thorough training, via the proposed routing masking method.
 Similarly to Share-MoE, MaskMoE also employs the remarkable shared expert architecture, which has been proven by  Dai et al.
@@
 (2023a, b); Huang et al.
-(2024), we adopt the LLaMA architecture with 24 transformer layers and a hidden layer dimension of 1024 and we employ the AdamW  Loshchilov and Hutter (2019) optimizer with a cosine learning rate decay schedule.
+(2024), we adopt the LLaMA architecture with 24 Transformer blocks and a hidden-state dimension of 1024.
+We employ the AdamW  Loshchilov and Hutter (2019) optimizer with a cosine learning rate decay schedule.
 For the MoE model, following  Roller et al.
 (2021), we conduct experiments under both single-layer and multi-layer settings.
-For the single-layer setting, we replace the final FFN sublayer with the MoE sublayer.
+For the single-layer setting, we replace the FFN layer with the MoE layer only for the last Transformer block.
 For the multi-layer MoE models, we follow Gshard Lepikhin et al.
-(2021), and the FFN sublayer with MoE sublayer for MoE layers every other layer, resulting in a total of $12$ MoE layers in this setting.
-For single-layer MoE models and dense models, we use a learning rate of $3e^{-4}$, while for $12$-layer MoE models, we employ a learning rate of $1e^{-4}$ to ensure stable convergence of the models.
-If not specified otherwise, tokens that account for the top 40% of frequency are considered frequent tokens, while the remaining are considered infrequent tokens, i.e., $P=0.4$.
+(2021), and replace the FFN layer with MoE layer for every other Transformer block, resulting in a total of $12$ MoE layers in this setting.
+For single-layer MoE models and dense models, we use a learning rate of $3e^{-4}$, while for Multi-layer MoE models, following  Lewis et al.
+(2021), we employ a learning rate of $1e^{-4}$ to ensure stable convergence of the models.
+If not specified otherwise, the top most frequent tokens that cover 40% of the training set are considered frequent tokens, while the remaining are considered infrequent tokens, i.e., $P=0.4$.
 Empirically, for frequent tokens, the visible expert count $V$ is set to 8, and for infrequent tokens, the visible expert count $V$ is set to 1.
@@
 (2021) implemented in the library to facilitate faster model convergence and better runtime efficiency.
-Following previous works Roller et al.
-(2021); Fedus et al.
-(2022); Dai et al.
-(2022), unless otherwise specified, our experiments only select the top expert.
 Model BoolQ Hellaswag LAMBADA PIQA SIQA StoryCloze Arc-e TriviaQA WebQs
@@
 (2024); Dai et al.
-(2024), we conduct tests on a large number of downstream benchmarks, including zero-shot tests for BoolQ Clark et al.
+(2024), we conduct tests on various downstream benchmarks, including zero-shot tests for BoolQ Clark et al.
 (2019), HellaSwag Zellers et al.
@@
 (2016), and Arc-e Bhakthavatsalam et al.
-(2021), as well as 5-shot tests for TriviaQA Joshi et al.
+(2021).
+Following  Touvron et al.
+(2023a); Su et al.
+(2024), we conduct 5-shot tests for TriviaQA Joshi et al.
 (2017) and WebQs Berant et al.
@@
 Table 1 shows the main results of language modeling on the Pile validation sets.
-With the same number of activated parameters during inference (i.e., 468M), MoE models consistently achieve improvements over the dense model.
-In comparison to SMoE and Hash Layer, under the Single-MoE-Layer setup, MaskMoE has respectively reduced the PPL on the Pile Validation by 0.14 and 0.08.
+With the same number of activated parameters during inference (i.e., 468M), MoE models consistently achieve improvements (i.e., substantially lower PPL) over the dense model.
+In comparison to SMoE and Hash Layer, under the Single-MoE-Layer setup, MaskMoE has respectively reduced the PPL on the Pile validation set by 0.14 and 0.08.
 In the Multi-MoE-Layer setup, their corresponding PPL are, respectively, reduced by 0.07 and 0.05.
-Compared to Share-MoE, MaskMoE achieves a PPL decrease of 0.24 in Single-MoE-Layer setup and another PPL decrease of 0.04 in Multi-MoE-Layer setup.
-Such results demonstrate the reasonableness of the MaskMoE design.
-It is noteworthy that Share-MoE achieves worse results than SMoE and Hash Layer in the Single-MoE-Layer setup.
-However, MaskMoE consistently maintains superior performance, whether in single or multi-MoE layers.
-It indicates that MaskMoE exhibits stronger robustness across various settings.
+Compared to Share-MoE, MaskMoE also achieves a PPL decrease of 0.24 in Single-MoE-Layer setup and 0.04 in Multi-MoE-Layer setup.
+Such results demonstrate the effectiveness of our proposed MaskMoE.
+It is noteworthy that Share-MoE achieves worse results than SMoE and Hash Layer in the Single-MoE-Layer setup, even with a built-in shared expert.
+However, MaskMoE consistently maintains superiority over them in both single-layer and multi-layer setups.
+That indicates MaskMoE exhibits stronger robustness across various settings.
 4.4.2 Benchmark Results
 As shown in Table 2, we present the performance of the models on downstream tasks.
-Firstly, we observe that MoE models consistently and significantly outperform the dense models in downstream tasks.
-And more importantly, we find that MaskMoE substantially outperforms the other MoE baselines, whether in Single-MoE-Layer setup or Multi-MoE-Layer setup.
-Compared to SMoE, MaskMoE consistently outperforms it in all 9 benchmarks with the Single-MoE-Layer setup and 8 out of 9 benchmarks with the Multi-MoE-Layer setup.
-Against Hash Layer, MaskMoE excels in 7 out of 9 benchmarks with the Single-MoE-Layer setup and in 8 out of 9 benchmarks with the Multi-MoE-Layer setup.
-Compared to Share-MoE, MaskMoE demonstrates significant improvements in 7 out of 9 benchmarks with the Single-MoE-Layer setup and in 6 out of 9 benchmarks with the Multi-MoE-Layer setup.
-The result above well demonstrates the performance enhancement achieved by using our proposed router masking method to control the visibility of experts for different tokens.
+We can observe that MoE models consistently and significantly outperform the dense models in downstream tasks.
+More importantly, MaskMoE substantially outperforms the other MoE baselines, whether in Single-MoE-Layer setup or Multi-MoE-Layer setup.
+Specifically, Compared to SMoE, MaskMoE consistently outperforms it in all 9 benchmarks with the Single-MoE-Layer setup and 8 of 9 benchmarks with the Multi-MoE-Layer setup.
+Against Hash Layer, MaskMoE excels in 7 of 9 benchmarks with the Single-MoE-Layer setup and in 8 of 9 benchmarks with the Multi-MoE-Layer setup.
+Compared to Share-MoE, MaskMoE demonstrates significant improvements in 7 of 9 benchmarks with the Single-MoE-Layer setup and in 6 of 9 benchmarks with the Multi-MoE-Layer setup.
+The result above well demonstrates the performance enhancement achieved by using our proposed router masking method that controls the visibility of experts for different tokens.
 We believe the improvements are due to two factors.
 Firstly, compared to SMoE and Share-MoE, infrequent tokens are only routed to a single expert, ensuring more thorough training for those tokens.
-Secondly, in contrast to the Hash Layer, frequent tokens have multiple expert options, which promotes diversity in the representation learning of frequent tokens.
+Secondly, in contrast to the Hash Layer, frequent tokens have multiple expert options, which promotes diversity in their representation learning.
 5 Analyses
@@
 To explore the impact of shared experts, we conduct ablation experiments on them (i.e., removing the shared experts), and the results are shown in Table 3.
-We separately report the average scores [^3] Due to the limited space, here we simply average all metrics on downstream tasks to get an average score for each model.
-of the benchmarks and the PPL of Pile Validation.
+We separately report the performance on the benchmark and the PPL on the Pile validation set.
 Note that after removing the shared experts, we adopt the same architecture as SMoE and Hash Layer, where the MoE layer consists of 64 fully-sized FFN.
-We find that even without shared experts, MaskMoE still outperforms SMoE, Hash Layer, and Share-MoE in average benchmark scores and PPL.
+We find that even without Shared-Experts, MaskMoE still outperforms SMoE, Hash Layer, and Share-MoE in terms of benchmark scores and PPL, with Table 2 as references.
+Specifically, compared to SMoE, MaskMoE w/o Shared-Experts excels in 8 of 9 benchmarks.
+Additionally, compared to Hash Layer and Share-MoE, MaskMoE w/o Shared-Experts gains significant improvements in 7 of 9 benchmarks.
+Similarly, MaskMoE w/o Shared-Experts, yields lower PPL than SMoE, Hash Layer, and Share-MoE, with Table 1 as references.
 Such a result further validates the soundness of our proposed MaskMoE.
-Moreover, incorporating shared expert structures continuously improves MaskMoE’s performance.
+Moreover, incorporating Shared-Expert structures continuously improves MaskMoE’s performance.
+As shown in Table 3, compared to MaskMoE w/o Shared-Expert, MaskMoE excels in 5 of 9 benchmarks and has a lower PPL in the Pile validation set.
 We attribute it to that the shared expert structure promotes greater specialization in token representation DeepSeek-AI et al.
 (2024).
-5.2 Impact of Router Mask
-Model Avg($\uparrow$) PPL($\downarrow$)
-SMoE 43.87 6.62
-Hash Layer 44.56 6.56
-Share-MoE 44.25 6.72
-MaskMoE 44.99 6.48
-w/o Shared Experts 44.85 6.51
-Table 3: Average scores benchmarks and PPL on the Pile validation sets. The best score is marked in bold, and the second best is underlined.
-Considering the proposed MaskMoE without shared experts still outperforms the compared MoE methods, to show the impact of router mask more significantly and minimize the impact of the shared expert, the experiments remove the shared expert.
-Instead, we adopt the same structure as SMoE and Hash Layer, with each MoE layer comprising 64 full-sized FNN modules as experts.
-The performance of MaskMoE is predominantly influenced by two hyperparameters: the boundary threshold for the separation of frequent and infrequent tokens, and the maximum number of visible experts (i.e., $V$) for tokens.
-As shown in Table 4, we conduct an experimental parameter search for both and report their PPL values on the Pile validation sets.
-$V_{a}$ represents the number of experts visible for frequent tokens, and $V_{b}$ represents the number of experts visible for infrequent tokens.
-Notably, when $V_{a}=V_{b}=1$, MaskMoE is equivalent to the Hash Layer, and when $V_{a}=V_{b}=64$, MaskMoE is equivalent to SMoE.
-Then we have the following observations.
-Firstly, we find that MaskMoE outperforms the Hash Layer.
-With $V_{a}=8$ and $V_{b}=1$, MaskMoE reduces ppl by 0.052 on the Pile validation set compared to the Hash Layer, i.e., $V_{a}=V_{b}=1$.
-The result indicates that the multiple-expert design in the proposed MaskMoE for frequent tokens does enhance their representation diversity, which is beneficial for model performance.
-Secondly, compared to SMoE, concentrating tokens into fewer visible experts, such as settings like $V_{a}=V_{b}=8$ or $V_{a}=V_{b}=4$, yields better outcomes.
-It indicates that routing tokens more densely to a reduced number of visible experts can enhance thorough training and improve model performance.
-Moreover, we observe that the configuration $V_{a}=8$ and $V_{b}=1$ outperforms $V_{a}=V_{b}=8$.
-It well verifies that further reducing the number of visible experts is more beneficial for infrequent tokens.
-Finally, it is noteworthy that for MaskMoE, when $V_{b}=1$, regardless of the value of $V_{a}$ (whether it is 4, 8, 16, or 64), the performance of MaskMoE consistently outperforms that of SMoE and Hash Layer.
-That also demonstrates the robustness of MaskMoE.
+MaskMoE w/o Shared-Expert
+BoolQ 58.38 57.40
+Hellaswag 44.47 44.10
+LAMBADA 55.36 54.86
+PIQA 68.99 69.10
+SIQA 41.86 41.97
+StoryCloze 65.15 65.69
+Arc-e 53.14 53.91
+TriviaQA 10.39 10.01
+WebQs 7.14 6.89
+PPL($\downarrow$) 6.48 6.51
+Table 3: Impact of Shared Expert on benchmark performances and the Pile validation PPL. The best score is marked in bold.
 $P$ $V_{a}$ $V_{b}$ PPL $P$ $V_{a}$ $V_{b}$ PPL
-0.2 8 1 6.518 0.4 16 1 6.509
-0.4 8 1 6.506 0.4 4 1 6.511
-0.6 8 1 6.518 0.4 64 1 6.549
-/ 64 64 6.618$\spadesuit$ / 1 1 6.558 $\clubsuit$
-/ 4 4 6.551 / 8 8 6.566
-Table 4: Impact of Router Mask. $\spadesuit$ denotes SMoE, $\clubsuit$ denotes Hash Layer.
-5.3 Number of Experts
-To investigate the variation in model performance under different configurations of expert numbers, we conducted experiments with 32, 64, and 128 experts.
-The experimental results, as shown in Figure 2, reveal the following observations.
-First of all, we observe that for SMoE and Share-MoE, increasing the number of experts does not necessarily lead to a lower PPL.
-In fact, when the number of experts reaches 128, both models even exhibit higher PPLs compared to having only 64 experts.
-We attribute it to that having more experts leads to the occurrences of an identical token being spread among more experts, resulting in each expert not learning enough about this token, which leads to model underfitting.
-In contrast, the performance of Hash Layer and MaskMoE continue to improve with the increase in the number of experts.
-We attribute it to that, despite the growing number of experts, the visible experts for an identical token in Hash Layer and MaskMoE remain a fixed, small number, allowing the experts to continue learning adequately.
-It indicates that they are more favorable for scaling the number of experts.
-Secondly, we observe that compared to the Hash Layer, MaskMoE consistently performs better across all configurations.
-Although the Hash Layer also enables experts to be adequately trained with the tokens assigned to them, MaskMoE excels in the diversity of representations learning for frequent tokens.
-such a superiority in handling token diversity contributes to MaskMoE’s overall advantage over Hash Layer.
-Finally, we find that regardless of the number of experts, MaskMoE consistently maintains the best performance.
-That further verifies the superiority of the proposed MaskMoE and the reasonableness of its design.
+0.0 8 1 6.558 $\spadesuit$ 0.6 8 1 6.518
+0.2 8 1 6.518 0.8 8 1 6.539
+0.4 8 1 6.506 1.0 8 1 6.566 $\spadesuit$
+Table 4: Impact of the frequency threshold $P$ for the separation of frequent and infrequent tokens. $\spadesuit$ denotes no distinction between frequent and infrequent tokens.
+$P$ $V_{a}$ $V_{b}$ PPL $P$ $V_{a}$ $V_{b}$ PPL
+/ 64 64 6.618$\spadesuit$ 0.4 64 1 6.549
+/ 8 8 6.566 0.4 8 1 6.506
+/ 4 4 6.551 0.4 4 1 6.511
+/ 1 1 6.558 $\clubsuit$ 0.4 16 1 6.509
+Table 5: Impact of visible experts. $\spadesuit$ denotes SMoE, $\clubsuit$ denotes Hash Layer.
+5.2 Impact of Hyperparameters for Router Masking
+Considering the proposed MaskMoE without shared experts still outperforms the compared MoE methods, to show the impact of router mask more clear and minimize the impact of the shared expert, here we keep the same experimental setting with the shared expert removed and conduct further experiments to analyze the impact of hyperparameters of our proposed MaskMoE.
+The performance of MaskMoE is predominantly influenced by two hyperparameters: the boundary threshold (i.e., $P$) for the separation of frequent and infrequent tokens, and the maximum number of visible experts (i.e., $V$) for tokens.
+As shown in Table 4 and 5, we conduct parameter searches for both and report their PPL values on the Pile validation sets.
+Specifically, here $V_{a}$ represents the number of experts visible for frequent tokens, and $V_{b}$ represents the number of experts visible for infrequent tokens.
+As shown in Table 4, we report the impact of the boundary threshold $P$.
+Here, $P=0.0,V_{a}=8,V_{b}=1$ and $P=1,V_{a}=8,V_{b}=1$ indicate no distinction between frequent and infrequent tokens, with every token having $1$ and $8$ visible experts, respectively.
+Their performances are significantly worse than those under the settings of $10<P<$ 1, given that $V_{a}=8,V_{b}=1$.
+That indicates tokens of different frequencies indeed require distinct visible experts, validating the effectiveness of our design.
+Additionally, we observe that regardless of the settings of $P$ where $10<P<$ 1, the PPLs are consistently lower than $P=0.0$ or $P=1$.
+That validates the robustness of MaskMoE.
+[…]
 (a) None-Shared Expert Structures
@@
 Figure 2: Comparison of MoE-based Transformers with different numbers of experts.
+5.3 Impact of the Number of Experts
+To investigate the variation in model performance under different configurations of expert numbers, we conducted experiments with 32, 64, and 128 experts.
+The experimental results are illustrated in Figure 2.
+First of all, we observe that for SMoE and Share-MoE, increasing the number of experts does not necessarily lead to a lower PPL.
+In fact, when the number of experts reaches 128, both models even exhibit higher PPLs compared to having only 64 experts.
+We attribute it to that having more experts leads to the occurrences of an identical token being spread among more experts, and that disables each expert to learn adequately about this token, leading to model underfitting.
+In contrast, the performance of Hash Layer and MaskMoE continue to improve as the number of experts increases.
+Actually, when the number of experts increases, the average number of token occurrences assigned to each expert decrease.
+However, Hash Layer and MaskMoE can still ensure that each expert is fed with all or a large portion of the occurrences of a token, though the number of distinct tokens decreases for each expert.
+That makes the data for each expert more consistent and easier to learn.
+Secondly, compared to Hash Layer, MaskMoE consistently performs better across all configurations.
+Although Hash Layer also enables experts to be adequately trained for specific tokens, MaskMoE excels in the diversity of representation learning for frequent tokens, which contributes to MaskMoE’s overall advantage over Hash Layer.
+Finally, regardless of the number of experts, MaskMoE consistently yields the best performance. That further verifies its superiority and reasonableness.
+5.4 Performance Analysis of Top-k Routing
+Model SMoE Hash Layer Share-MoE MaskMoE
+PPL($\downarrow$) 6.55 6.50 6.58 6.46
+Table 6: The PPL on the Pile validation sets under the top-$2$ gating. The best score is marked in bold.
+We further analyze the performance of the MoE models under the top-$k$ ($k=2$) routing mechanism.
+In the SMoE model, the top-$2$ experts are selected from 64 fully-sized FFN experts based on the router’s output scores.
+For the Hash Layer, two experts are randomly chosen from 64 fully-sized FFN experts using a random hash function.
+Regarding Share-MoE and MaskMoE models, in addition to the shared experts, the top-$2$ experts are selected from the 128 half-sized FFN experts based on the router’s output scores.
+All settings share the same number of active parameters.
+Note that in MaskMoE, the number of visible experts is doubled from the original setting, i.e., $V_{a}=16,V_{b}=2$.
+As shown in Table 6, the model’s PPLs on the Pile validation sets are presented.
+MaskMoE consistently yields the best performance, achieving a 0.12 decrease in PPL compared to Share-MoE.
+That further verifies the effectiveness.
 6 Conclusions
-In existing MoE language models, the widely-used dynamic routing models such as SMoE, which routes the same token to multiple experts, potentially cause the model to encounter underfitting during training, especially for infrequent tokens.
+In existing MoE language models, the widely-used dynamic routing methods such as SMoE, exhibit routing fluctuations, and route the occurrences of a token to multiple experts during training, potentially causing the model to encounter underfitting, especially for infrequent tokens.
 Although fixed routing like Hash Layer can mitigate the routing fluctuation issue, it struggles with capturing rich features for frequent tokens.
-To tackle the routine fluctuation issue and the representation diversity learning issue at the meantime, we propose a routing masking method termed MaskMoE, aiming at ensuring adequate learning of infrequent tokens while maintaining rich representations learning for frequent tokens.
-Extensive experiments demonstrate that the proposed MaskMoE outperforms existing MoE methods like SMoE, Hash Layer, and Share-MoE on various downstream tasks and significantly reduces PPL on the Pile validation sets.
-That will demonstrates the effectiveness of MaskMoE.
+Considering both routing fluctuation and representation diversity, we propose MaskMoE, a routing masking method that ensures adequate learning for infrequent tokens and rich representations for frequent tokens.
+Extensive experiments show that MaskMoE outperforms SMoE, Hash Layer, and Share-MoE on various downstream tasks and significantly reduces PPL on the Pile validation sets.
 7 Limitiations